### v1

# Large Language Models for Secure Code Assessment: A Multi-Language Empirical Study

Abstract
Most vulnerability detection studies focus on datasets of vulnerabilities in C/C++ code, offering limited language diversity. Thus, the effectiveness of deep learning methods, including large language models (LLMs), in detecting software vulnerabilities beyond these languages is still largely unexplored.
In this paper, we evaluate the effectiveness of LLMs in detecting and classifying Common Weakness Enumerations (CWE) using different prompt and role strategies. Our experimental study targets six state-of-the-art pre-trained LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLLama-7B, CodeLLama-13B, and Gemini 1.5 Pro) and five programming languages: Python, C, C++, Java, and JavaScript. We compiled a multi-language vulnerability dataset from different sources, to ensure representativeness. Our results showed that GPT-4o achieves the highest vulnerability detection and CWE classification scores using a few-shot setting.
Aside from the quantitative results of our study, we developed a library called CodeGuardian integrated with VSCode which enables developers to perform LLM-assisted real-time vulnerability analysis in real-world security scenarios. We have evaluated CodeGuardian with a user study involving 22 developers from the industry. Our study showed that, by using CodeGuardian, developers are more accurate and faster at detecting vulnerabilities.
I Introduction
A software vulnerability is a defect that could allow an attacker to gain control of a software system, steal or manipulate sensitive data, install a backdoor, or plant other types of malware [1, 2, 3]. Software vulnerabilities vary in nature and level of exploitability. Not all vulnerabilities allow attackers to cause the same types of harm, and not all vulnerabilities create equally severe risks. However, all vulnerabilities pose at least some level of risk to the applications they impact, as well as the environments that host those applications.
Software vulnerabilities may have different root causes, spanning from the way the software is designed, problems with the software’s source code, poor management of data, or access control settings within the application. In this work, we focus on classes of vulnerabilities related to bugs in the source code. The U.S. Department of Homeland Security reports that 90% of security incidents are caused by exploiting security flaws and defects in the code [4]. A recent survey [5] found that over 55% of developers struggle to identify code vulnerabilities, indicating a lack of secure coding knowledge. Due to the increasing frequency and severity of cybersecurity incidents, which negatively impact end-users, businesses, and critical infrastructure [6], identifying security vulnerabilities within source code has become essential.
Researchers have proposed methods for the automatic detection and repair of software vulnerabilities [7, 8, 9, 10, 11, 12, 13]. For instance, program analysis-based techniques have been widely employed [14]. However, such rule-based techniques are affected by high false positive rates and low effectiveness in handling a large number of diverse types of vulnerabilities [12]. In addition, they often suffer from long response times [13], which makes them less suitable for continuous integration/development environments. Recently, Deep Learning (DL) techniques have enabled significant progress in vulnerability detection in the source code [15, 16, 17]. However, most DL models treat the code as linear sequences or textual data, which limits the finding of the semantic relationship within the source code and demands the development of more robust and efficient methods.
To address these limitations, researchers have started exploring the potential of Large Language Models (LLMs) that are pre-trained on a large corpus of data and use advanced natural language processing capabilities based on the transformer architecture [18]. LLMs have shown great results across various natural language and software engineering tasks [18, 11], including automated vulnerability detection and repair [9, 10, 11]. Existing studies on vulnerability detection are executed against datasets that include vulnerabilities pertaining to only one or two programming languages. Instances are the vulnerability-fixing commit dataset by Zhou et al. [19] (C/C++), by Li et al. [20] (Java), BigVul [21] (C/C++), SVEN [22] (Python and C/C++), Devign [23] (C), D2A [24] (C/C++). Consequently, the capabilities of LLMs across diverse programming languages remain largely unexplored. Furthermore, these studies are purely quantitative, and the potential of LLMs to assist developers in identifying vulnerabilities has not yet been investigated.
To fill these research gaps, this paper investigates the effectiveness of LLMs in identifying vulnerable codes across different programming languages. To the best of our knowledge, this is the first study that reports an extensive evaluation of six state-of-the-art pre-trained LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLLama-7B, CodeLLama-13B, and Gemini 1.5 Pro) in identifying the vulnerabilities pertaining to five programming languages, namely Python, C, C++, Java, and JavaScript. Our study focuses on the top 25 Common Weakness Enumeration (CWE) classes that represent the most dangerous software weaknesses [25], for which we compiled a dataset of vulnerable snippets from existing reference datasets. Our study targets vulnerability detection and CWE classification using zero-shot and few-shot approaches.
Furthermore, to investigate the industrial applicability of LLMs in software vulnerability detection and debugging, we have developed a VSCode extension called CodeGuardian. It integrates the LLM analyzed by our quantitative study to allow developers to scan code dynamically while editing, providing instant feedback and enhancing secure coding practices directly within their workflow. We have evaluated CodeGuardian with a user study involving six tasks across the five programming languages and 22 participants with software engineering experience in the industry and varying levels of expertise in software security. Our qualitative analysis revealed that by using CodeGuardian, developers were twice as accurate at CWE classification and 60% faster, thanks to the LLM-powered engine.
Our paper makes the following contributions:
Dataset. A dataset of more than 370 manually validated vulnerabilities, pertaining to five programming languages, which can support the evaluation of automated techniques in security vulnerability detection and classification.
Evaluation. An empirical study concerning the capabilities of six pre-trained LLMs at detecting and classifying software vulnerabilities in five programming languages.
CodeGuardian. A VSCode extension that allows developers to use LLMs to scan code for vulnerabilities while editing.
User Study. A qualitative analysis of CodeGuardian with a user study of 22 developers from our industrial partner.
II Empirical Study
II-A Research Questions
Our study considers the following research questions:
RQ1 (vulnerability detection): How effective are LLMs in detecting CWEs across multiple languages? How does effectiveness vary among different languages?
RQ2 (CWE classification): How effective are LLMs in classifying CWEs across multiple languages? How does effective vary across languages? How does effectiveness vary among different languages?
The first research question evaluates how accurately LLMs perform binary classification, determining whether a code snippet is vulnerable or not. The second research question evaluates the accuracy of LLMs in classifying CWEs by assigning the correct CWE ID to a given code snippet.
II-B Dataset Design
The representativeness of datasets is crucial for effective vulnerability detection and classification [26, 27]. Existing works adopted datasets with limited language diversity, focusing primarily on C/C++ [19, 20, 21, 22, 23, 24].
In contrast, in this work, we consider software vulnerabilities of five popular programming languages, namely Python, C, C++, Java, and JavaScript. We compiled a dataset that integrates three existing vulnerability datasets: CVEFixes, CWE-snippets, and the JavaScript Vulnerability DataSet (JVD). Table I shows the statistics of these datasets, including the number of languages, CWEs, and files. CVEFixes is a fix-commit vulnerability dataset [28]. CWE-snippets is a dataset created in collaboration with an industrial partner for this study. It consists of numerous vulnerability snippets across multiple programming languages. JVD is a vulnerability dataset for JavaScript code [29]. From each dataset, we selected the top 25 CWE most dangerous software weaknesses [25] for each of the five selected programming languages.
Table I: Vulnerabilities benchmarks evaluated in this study.
      Dataset       # Languages       # CWEs       # Files
      CVEFixes [28]       27       180       18,249
      CWE-snippets       42       506       231,571
      JVD [29]       1       N/A       12,125
The maximum number of snippets per class was set to three across all the languages due to a limited number of vulnerable code samples available in the reference datasets for certain CWE classes and programming languages. For example, CWE-787 (Out-of-bounds Write) is more commonly associated with memory-unsafe languages like C and C++, while it is less prevalent in memory-safe languages such as Python and Java. Consequently, we prioritized constructing a balanced representation of vulnerabilities across all languages in our dataset. As a result, we aimed to construct a balanced representation of vulnerabilities across all languages in our dataset. In total, we retained 378 snippets, ensuring an equal number of vulnerable and non-vulnerable snippets for each programming language ($n_{Python}=38$, $n_{Java}=42$, $n_{C++}=36$, $n_{C}=44$, and $n_{JavaScript}=29$).
Table II: LLMs evaluated for vulnerability detection and classification.
Model Parameters Architecture Input Tokens Output Tokens
GPT-3.5-Turbo [30] $>$ 100B Decoder 16,385 4,096
GPT-4 Turbo [31] $>$ 100B Decoder 128K 4,096
GPT-4o [32] $>$ 100B Decoder 128K 4,096
CodeLLama-7B [33] 7 B Decoder 100K 4,096
CodeLLama-13B [33] 13 B Decoder 100K 4,096
Gemini 1.5 Pro [34] $>$ 100B Decoder 1M 8,192
Table III: System and user prompts: VD-Sys1, VD-Sys2, VD-User1 and VD-User2 are used for vulnerability detection (RQ1). CWE-Sys1 and CWE-Sys2 are used for system prompts 1 and 2 for CWE Classification. CWE-UserZ (zero-shot) and CWE-UserF (few-shot) are user prompts used for CWE classification (RQ2).
No. Prompt Type Prompt Template Verbalizer
VD-Sys1 System Description You are an AI binary vulnerability classifier that identifies whether the provided code is vulnerable or not vulnerable. You should respond with either only ’vulnerable’ or ’not vulnerable’. N/A
VD-Sys2 System Description You are an experienced developer who knows the security vulnerability very well. N/A
VD-User1 User Description for VD Classify the following code in vulnerable or not vulnerable. Output either only ’vulnerable’ or ’not vulnerable’. +: vulnerable
-: non-vulnerable
VD-User2 User Description for VD Now you need to identify whether a code contains a vulnerability or not. If has any potential vulnerability, output: ’vulnerable’. Otherwise, output: ’not vulnerable’.
You must respond with either ’vulnerable’ or ’not vulnerable’ only. The code is below.
+: vulnerable
-: non-vulnerable
CWE-Sys1 System Description for CWE You are an AI vulnerability classifier that identifies CWE of a sent code, if that code has a vulnerability.
You should respond with only one CWE. You don’t have to respond with any other information except for one of the CWEs.
N/A
CWE-Sys2 System Description for CWE Classification You are an experienced developer who knows the security vulnerability very well. N/A
CWE-UserZ User Description for CWE (zero-shot) Classify the following code in CWE a category. Output only one CWE tag in lowercase letters. Do not provide the full name of the CWE. Respond with ’non-vul’ if you think the code is not vulnerable. CWE ID
CWE-UserF User Description for CWE Classification with the few-shot setting Classify the following code in a CWE category. The code should have a vulnerability that corresponds to one of the top-25 CWEs below. Output only one CWE tag in lowercase letters. Do not provide the full name of the CWE. Respond ’non-vul’ if you think the code is not vulnerable. Below is the list of the top 25 CWEs:
CWE-787: Out-of-bounds Write
CWE-276: Incorrect Default Permissions.
CWE ID
II-C Large Language Models
We evaluate the effectiveness of using six state-of-the-art pre-trained LLMs. Table II shows the selected LLMs, including their parameter size (the number of parameters for GPT and Gemini 1.5 Pro models are not disclosed by the organizations, and therefore, we report an estimate), architecture, input, and output token sizes. Specifically, we considered GPT-3.5 Turbo, GPT-4 Turbo, and CodeLlama models available as of June 2024, and pertaining from different organizations such as OpenAI’s GPT-3.5-Turbo [30], GPT-4 Turbo [31] and GPT-4o [32], as well as Meta’s CodeLLama-7B [33] and CodeLLama-13B [33], and Google’s Gemini 1.5 Pro [34].
For the GPT models, we utilized the supporting company’s Azure OpenAI Studio. Concerning hosting CodeLLama-7B and CodeLLama-13B models, we employed Google Colab and the transformers library by HuggingFace [35]. For the Gemini 1.5 Pro model, we used the Vertex AI from the Google Cloud Platform. Lastly, to minimize the models’ creativity and obtain more focused, conservative, and consistent responses, we set the temperature parameter to $0.1$ for all LLMs, as done in previous studies [36].
Table IV: RQ1: Effectiveness comparison of six LLMs using four prompt configurations for CWE detection. The best scores are indicated in bold.
LLM / Prompt VD-Sys1 + VD-User1 VD-Sys1 + VD-User2 VD-Sys2 + VD-User1 VD-Sys2 + VD-User2
Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$
All
GPT-3.5 Turbo 0.62 0.59 0.77 0.67 0.57 0.55 0.82 0.66 0.72 0.73 0.70 0.71 0.61 0.58 0.77 0.66
GPT-4 Turbo 0.81 0.81 0.82 0.81 0.79 0.73 0.92 0.81 0.79 0.75 0.87 0.81 0.78 0.71 0.93 0.81
GPT-4o 0.80 0.72 0.97 0.83 0.71 0.64 0.99 0.78 0.62 0.57 1.00 0.72 0.63 0.57 0.99 0.73
CodeLlama-7b 0.62 0.58 0.89 0.70 0.57 0.54 0.93 0.68 0.54 0.52 0.98 0.68 0.57 0.55 0.84 0.66
CodeLlama-13b 0.60 0.76 0.28 0.41 0.66 0.67 0.65 0.66 0.65 0.62 0.77 0.69 0.66 0.62 0.81 0.71
Gemini 1.5 Pro 0.80 0.79 0.83 0.81 0.69 0.65 0.84 0.73 0.75 0.71 0.83 0.77 0.68 0.62 0.97 0.75
Python
GPT-3.5 Turbo 0.63 0.59 0.87 0.70 0.57 0.54 0.89 0.67 0.74 0.75 0.71 0.73 0.61 0.57 0.87 0.69
GPT-4 Turbo 0.75 0.73 0.79 0.76 0.78 0.72 0.89 0.80 0.78 0.73 0.87 0.80 0.71 0.66 0.87 0.75
GPT-4o 0.76 0.70 0.92 0.80 0.66 0.60 0.95 0.73 0.58 0.54 1.00 0.70 0.59 0.55 0.97 0.70
CodeLlama-7b 0.70 0.67 0.76 0.72 0.58 0.56 0.79 0.65 0.54 0.52 0.97 0.68 0.57 0.55 0.76 0.64
CodeLlama-13b 0.57 0.69 0.24 0.35 0.59 0.59 0.61 0.60 0.59 0.59 0.63 0.61 0.54 0.53 0.63 0.58
Gemini 1.5 Pro 0.76 0.79 0.71 0.75 0.61 0.59 0.71 0.64 0.74 0.70 0.84 0.76 0.66 0.60 0.95 0.73
C
GPT-3.5 Turbo 0.56 0.55 0.68 0.61 0.47 0.48 0.70 0.57 0.69 0.68 0.73 0.70 0.57 0.56 0.68 0.61
GPT-4 Turbo 0.83 0.78 0.91 0.84 0.75 0.67 0.98 0.80 0.80 0.75 0.89 0.81 0.78 0.70 0.98 0.82
GPT-4o 0.83 0.75 1.00 0.85 0.70 0.63 1.00 0.77 0.64 0.58 1.00 0.73 0.66 0.59 1.00 0.75
CodeLlama-7b 0.59 0.55 0.93 0.69 0.59 0.55 0.98 0.70 0.52 0.51 0.95 0.67 0.52 0.51 0.86 0.64
CodeLlama-13b 0.65 0.93 0.32 0.47 0.69 0.67 0.77 0.72 0.65 0.60 0.86 0.71 0.77 0.72 0.89 0.80
Gemini 1.5 Pro 0.77 0.73 0.86 0.79 0.70 0.67 0.80 0.73 0.78 0.79 0.77 0.78 0.73 0.65 1.00 0.79
C++
GPT-3.5 Turbo 0.60 0.57 0.75 0.65 0.60 0.56 0.86 0.68 0.72 0.71 0.75 0.73 0.61 0.59 0.75 0.66
GPT-4 Turbo 0.85 0.82 0.89 0.85 0.83 0.76 0.97 0.85 0.78 0.71 0.94 0.81 0.81 0.73 0.97 0.83
GPT-4o 0.86 0.78 1.00 0.88 0.75 0.67 1.00 0.80 0.62 0.57 1.00 0.73 0.60 0.55 1.00 0.71
CodeLlama-7b 0.61 0.57 0.92 0.70 0.57 0.54 0.94 0.69 0.57 0.54 1.00 0.70 0.60 0.56 0.94 0.70
CodeLlama-13b 0.61 0.83 0.28 0.42 0.62 0.63 0.61 0.62 0.68 0.64 0.83 0.72 0.69 0.64 0.89 0.74
Gemini 1.5 Pro 0.88 0.85 0.92 0.88 0.72 0.67 0.89 0.76 0.72 0.67 0.89 0.76 0.68 0.61 0.97 0.75
JavaScript
GPT-3.5 Turbo 0.74 0.69 0.86 0.77 0.69 0.63 0.93 0.75 0.81 0.85 0.76 0.80 0.72 0.67 0.90 0.76
GPT-4 Turbo 0.84 0.88 0.79 0.84 0.83 0.77 0.93 0.84 0.78 0.74 0.86 0.79 0.72 0.67 0.90 0.76
GPT-4o 0.67 0.61 0.97 0.75 0.66 0.59 1.00 0.74 0.53 0.52 1.00 0.68 0.53 0.52 0.97 0.67
CodeLlama-7b 0.64 0.58 0.97 0.73 0.55 0.53 0.97 0.68 0.52 0.51 1.00 0.67 0.66 0.60 0.90 0.72
CodeLlama-13b 0.53 0.57 0.28 0.37 0.69 0.72 0.62 0.67 0.66 0.64 0.72 0.68 0.62 0.58 0.86 0.69
Gemini 1.5 Pro 0.88 0.87 0.90 0.88 0.64 0.60 0.86 0.70 0.71 0.67 0.83 0.74 0.60 0.56 1.00 0.72
Java
GPT-3.5 Turbo 0.60 0.57 0.74 0.65 0.58 0.56 0.76 0.65 0.65 0.69 0.57 0.62 0.58 0.57 0.69 0.62
GPT-4 Turbo 0.80 0.86 0.71 0.78 0.79 0.76 0.83 0.80 0.82 0.83 0.81 0.82 0.83 0.78 0.93 0.85
GPT-4o 0.82 0.75 0.98 0.85 0.79 0.70 1.00 0.82 0.68 0.61 1.00 0.76 0.71 0.64 1.00 0.78
CodeLlama-7b 0.56 0.54 0.88 0.67 0.54 0.52 0.95 0.67 0.54 0.52 1.00 0.68 0.56 0.54 0.76 0.63
CodeLlama-13b 0.60 0.75 0.29 0.41 0.71 0.78 0.60 0.68 0.68 0.65 0.79 0.71 0.65 0.62 0.79 0.69
Gemini 1.5 Pro 0.75 0.74 0.76 0.75 0.76 0.70 0.93 0.80 0.77 0.74 0.83 0.79 0.71 0.65 0.95 0.77
II-D Classification Methods
Our study examined three classification approaches:
  • Vulnerability Detection: A simple binary classification method in which the LLM is tasked to categorize an input snippet into either “vulnerable” or “not vulnerable”.
  • CWE classification (zero-shot): A multi-class classification method in which the LLM is tasked to categorize an input snippet as pertaining to a specific CWE. The zero-shot setting provides only a task description in natural language as part of the input prompt.
  • CWE classification (few-shot): Same multi-class classification method as before, but using a few-shot setting in which the LLM is provided with code examples from the list of the top 25 CWEs [25].
II-E Prompt and Input Data
LLM’s prompts were divided into two main types: system prompt, which defines the behavior, and user prompt, which gives instructions regarding a given task.
In our study, the final prompts combine one system prompt with one user prompt. Table III reports all the prompts that were used for our experiments. We considered two system prompts: VD-Sys1 gives the LLM a detailed role for vulnerability detection. In contrast, VD-Sys2 had a more generic role, as described by Zhou et al. [37]. Similarly, we designed two different types of user prompts to detect a vulnerability: VD-User1 presents a concise request for vulnerability detection, whereas VD-User2 provides more descriptive instruction as used by Zhou et al. [37].
For the classification of CWE (RQ2), we introduced two system prompts (CWE-Sys1 and CWE-Sys2) and two user prompts (CWE-UserZ and CWE-UserF).
CWE-Sys1 is a similar prompt compared to VD-Sys1, which provides a more specific role for CWE classification. Conversely, CWE-Sys2 is the same prompt as VD-Sys2 [37].
CWE-UserZ is a user prompt for CWE classification with a zero-shot setting (i.e., without any examples). CWE-UserF is for CWE classification experiments with the few-shot setting, and it has a list of the top 25 CWEs comparable to CWE-UserZ. To form the final prompt text, we concatenated an input code snippet with one of our described prompt strategies.
II-F Procedure and Metrics
To answer RQ1 and RQ2, we executed each LLM on all code snippets of our datasets, using the prompt strategies described in Section II-E. As evaluation metrics, we employed four metrics, namely accuracy, recall, precision, and $F_{1}$ score, commonly used to evaluate the performance of LLMs in vulnerability detection.
In our evaluation, valid detections/classifications were treated as the positive class, while detection/classification augmentations were considered the negative class. We accounted for CWE hierarchical relationships, considering positive scores for cases where the LLM predicted a code snippet labeled with a specific CWE (e.g., CWE-77) as one of its child CWEs (e.g., CWE-78), according to the MITRE CWE database [38]. For CWE classification (RQ2), we evaluated the overall performance by macro-averaging the precision parameter across all classes in the precision, recall, and $F_{1}$ score.
Overall, in our evaluation, we computed 18,144 predictions (6 LLMs $\times$ 378 code snippets $\times$ 8 prompt configurations), approximately 10 hours of computing time.
Table V: RQ2: Effectiveness comparison of six LLMs using four prompt configurations for CWE classification. The best scores are indicated in bold.
LLM / Prompt CWE-Sys1 + CWE-UserZ CWE-Sys2 + CWE-UserZ CWE-Sys1 + CWE-UserF CWE-Sys2 + CWE-UserF
Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$
All
GPT-3.5 Turbo 0.45 0.14 0.12 0.12 0.38 0.14 0.12 0.12 0.52 0.34 0.35 0.32 0.44 0.24 0.23 0.21
GPT-4 Turbo 0.56 0.21 0.18 0.18 0.67 0.30 0.23 0.25 0.61 0.42 0.52 0.44 0.69 0.54 0.58 0.53
GPT-4o 0.62 0.23 0.19 0.20 0.52 0.22 0.19 0.19 0.67 0.61 0.69 0.61 0.65 0.60 0.74 0.63
CodeLlama-7b 0.22 0.12 0.09 0.08 0.29 0.08 0.05 0.06 0.11 0.13 0.10 0.08 0.13 0.11 0.08 0.06
CodeLlama-13b 0.33 0.09 0.08 0.08 0.25 0.08 0.05 0.05 0.23 0.09 0.07 0.07 0.21 0.10 0.09 0.07
Gemini 1.5 Pro 0.53 0.15 0.13 0.13 0.55 0.16 0.15 0.14 0.66 0.52 0.56 0.51 0.69 0.56 0.52 0.49
Python
GPT-3.5 Turbo 0.39 0.25 0.21 0.22 0.38 0.22 0.20 0.20 0.47 0.27 0.32 0.28 0.46 0.32 0.30 0.29
GPT-4 Turbo 0.61 0.43 0.42 0.40 0.66 0.42 0.35 0.36 0.51 0.53 0.54 0.48 0.59 0.59 0.55 0.52
GPT-4o 0.54 0.36 0.26 0.27 0.51 0.35 0.31 0.30 0.53 0.58 0.52 0.51 0.58 0.60 0.65 0.60
CodeLlama-7b 0.13 0.09 0.07 0.07 0.26 0.16 0.10 0.11 0.01 0.00 0.02 0.01 0.11 0.15 0.14 0.11
CodeLlama-13b 0.29 0.16 0.15 0.14 0.24 0.10 0.09 0.08 0.21 0.17 0.13 0.11 0.18 0.17 0.18 0.12
Gemini 1.5 Pro 0.54 0.24 0.21 0.20 0.50 0.25 0.22 0.22 0.67 0.61 0.64 0.58 0.71 0.72 0.62 0.62
C
GPT-3.5 Turbo 0.43 0.21 0.20 0.19 0.34 0.20 0.19 0.19 0.44 0.46 0.43 0.39 0.41 0.32 0.33 0.31
GPT-4 Turbo 0.60 0.38 0.37 0.36 0.70 0.46 0.42 0.43 0.65 0.56 0.66 0.58 0.70 0.67 0.69 0.66
GPT-4o 0.61 0.37 0.31 0.32 0.48 0.33 0.28 0.29 0.66 0.64 0.74 0.65 0.60 0.62 0.76 0.64
CodeLlama-7b 0.22 0.19 0.15 0.15 0.19 0.08 0.06 0.07 0.16 0.12 0.20 0.13 0.15 0.12 0.17 0.12
CodeLlama-13b 0.31 0.18 0.15 0.15 0.22 0.15 0.12 0.12 0.22 0.14 0.13 0.12 0.24 0.20 0.21 0.18
Gemini 1.5 Pro 0.40 0.22 0.17 0.18 0.53 0.29 0.26 0.26 0.67 0.57 0.60 0.56 0.70 0.67 0.62 0.61
C++
GPT-3.5 Turbo 0.43 0.21 0.19 0.19 0.36 0.25 0.22 0.22 0.54 0.43 0.46 0.41 0.43 0.32 0.32 0.29
GPT-4 Turbo 0.53 0.39 0.39 0.37 0.62 0.44 0.42 0.41 0.64 0.59 0.68 0.60 0.68 0.57 0.66 0.57
GPT-4o 0.65 0.42 0.39 0.39 0.51 0.40 0.38 0.36 0.65 0.61 0.75 0.64 0.60 0.59 0.76 0.62
CodeLlama-7b 0.33 0.17 0.15 0.14 0.36 0.18 0.12 0.14 0.15 0.17 0.18 0.16 0.11 0.09 0.13 0.09
CodeLlama-13b 0.29 0.18 0.15 0.15 0.25 0.12 0.11 0.10 0.26 0.22 0.20 0.19 0.24 0.27 0.22 0.19
Gemini 1.5 Pro 0.58 0.32 0.28 0.29 0.62 0.32 0.30 0.30 0.60 0.45 0.40 0.39 0.64 0.45 0.42 0.41
JavaScript
GPT-3.5 Turbo 0.53 0.25 0.26 0.24 0.47 0.27 0.27 0.25 0.64 0.42 0.50 0.44 0.50 0.41 0.40 0.37
GPT-4 Turbo 0.52 0.28 0.28 0.26 0.76 0.52 0.52 0.50 0.53 0.52 0.64 0.53 0.79 0.68 0.71 0.66
GPT-4o 0.69 0.37 0.42 0.38 0.53 0.29 0.29 0.27 0.74 0.59 0.62 0.59 0.69 0.53 0.61 0.55
CodeLlama-7b 0.21 0.16 0.20 0.17 0.36 0.25 0.25 0.22 0.14 0.13 0.19 0.13 0.14 0.05 0.12 0.06
CodeLlama-13b 0.41 0.26 0.30 0.26 0.38 0.27 0.27 0.23 0.17 0.08 0.07 0.06 0.14 0.13 0.11 0.09
Gemini 1.5 Pro 0.66 0.31 0.35 0.33 0.59 0.27 0.28 0.26 0.72 0.60 0.73 0.63 0.76 0.52 0.52 0.49
Java
GPT-3.5 Turbo 0.46 0.19 0.16 0.16 0.39 0.22 0.18 0.19 0.55 0.28 0.31 0.29 0.42 0.22 0.23 0.21
GPT-4 Turbo 0.51 0.29 0.23 0.25 0.63 0.42 0.35 0.37 0.68 0.49 0.55 0.50 0.69 0.48 0.52 0.48
GPT-4o 0.62 0.35 0.30 0.31 0.58 0.36 0.32 0.32 0.76 0.59 0.63 0.59 0.77 0.63 0.68 0.64
CodeLlama-7b 0.21 0.14 0.10 0.10 0.31 0.11 0.11 0.10 0.10 0.07 0.11 0.07 0.14 0.09 0.13 0.08
CodeLlama-13b 0.36 0.17 0.18 0.17 0.23 0.16 0.16 0.14 0.29 0.14 0.15 0.12 0.24 0.15 0.12 0.12
Gemini 1.5 Pro 0.51 0.28 0.26 0.26 0.51 0.29 0.26 0.26 0.65 0.59 0.69 0.59 0.65 0.52 0.60 0.52
III Experimental Results
III-A RQ1 (Vulnerability detection)
Table IV (vulnerability detection) presents the effectiveness results across the different prompt setups. Overall, GPT-4 Turbo and GPT-4o showed the best binary classification scores. The former achieved the highest accuracy and precision (0.81), whereas the latter showed the highest recall (1.00) and $F_{1}$ score (0.83). Changing the prompt from VD-Sys1 to VD-Sys2 (i.e., a more general role definition) had a positive impact only for GPT-3.5 Turbo and CodeLlama-13b, proving detrimental for the other LLMs. Although VD-Sys2 configurations achieved the highest recall results, they only marginally outperformed VD-Sys1 configurations and came at the cost of reduced precision. The best balance between detecting vulnerabilities and minimizing false positives was achieved by the prompt configurations VD-Sys1 + VD-User1 and VD-Sys1 + VD-User2 (the first two macro-columns). Among these, VD-Sys1 + VD-User1 was the most effective for minimizing false alarms (GPT-4 Turbo) and detecting security vulnerabilities (GPT-4o).
For Python, GPT-4 Turbo and GPT-4o consistently outperformed other models. In particular, both models achieved the highest $F_{1}$ score of 0.80 using the VD-User1 task prompt. Although Gemini 1.5 pro delivered the highest accuracy, its $F_{1}$ score did not reach more than 0.80 due to its lower recall than GPT-4 Turbo and GPT-4o across all the prompt settings.
In C, GPT-4o demonstrated exceptional recall, achieving perfect scores (1.00) across all prompt configurations. Although the model size of CodeLlama-13b is considerably smaller than other bigger models, such as GPT-4o and GPT-4 Turbo, it achieved the highest precision of 0.93. This revealed the potential to reliably identify true positive vulnerabilities with a smaller sized LLM.
For C++, both GPT-4 Turbo and Gemini 1.5 Pro showed a solid performance, both achieving the highest $F_{1}$ score of 0.88 using VD-Sys1 + VD-User1. Particularly, GPT-4o surpassed other models in recall (1.0) across all prompt sets, which is similar to C.
In JavaScript, Gemini 1.5 Pro maintained a balance between precision (0.87) and recall (0.90) in VD-Sys1 + VD-User1; thus, it achieved the highest $F_{1}$ score (0.88). Although GPT-4o obtained the highest recall of 1.0 using two prompt sets, CodeLlama-7b, a model more than 10 times smaller in size, also reached a recall of 1.0 (VD-Sys2 + VD-User1).
Lastly, regarding the result in Java, GPT-4 Turbo and GPT-4o showed the best $F_{1}$ score of 0.85 using VD-Sys1 + VD-User1 used by GPt-4o and VD-Sys2 + VD-User2 used by GPT-4 Turbo. While other models did not achieved satisfactory performance, GPT-4 Turbo exhibited the highest precision across all prompt sets. In particular, VD-Sys1 + VD-User1, the highest precision of 0.86 (GPT-4 Turbo), was approximately 15% better than the second highest one (GPT-4o).
RQ1 (vulnerability detection): Overall, GPT-4 Turbo and GPT-4o achieved the best vulnerability detection scores across multiple languages ($F_{1}=0.81$ and $F_{1}=0.83$). GPT-4 Turbo was the most effective for minimizing false alarms. GPT-4 Turbo achieved the highest $F_{1}$ score in Python. GPT-4o demonstrated exceptional recall, particularly in C and C++. Gemini 1.5 Pro performed the best in JavaScript. GPT-4-based models were the most effective LLMs in Java.
III-B RQ2 (CWE classification)
Table V (CWE classification) presents the effectiveness results across the different prompt setups. Unlike the detection task, the scores vary significantly across configurations. Similarly to the detection task, overall, GPT-4 Turbo and GPT-4o showed the best multi-class classification scores. Classification demonstrates a more difficult task than classification, as GPT-4 Turbo achieved an $F_{1}$ score of $0.53$ and GPT-4o achieved an $F_{1}$ score of $0.63$. The best results are obtained using the few-shot learning technique and a generic role (CWE-Sys2 + CWE-UserF configuration).
The effectiveness of GPT-3.5 Turbo, GPT-4 Turbo, GPT-4o, and Gemini 1.5 Pro, improved using the few-shot learning technique. GPT-4o’s $F_{1}$ score improves by approximately tripling in all prompt sets when switching from zero-shot to few-shot. Similarly, GPT-4 Turbo shows substantial improvements with more than double the recall and $F_{1}$ score between CWE-Sys2 + CWE-UserZ and CWE-Sys2 + CWE-UserF.
For Python, Gemini 1.5 Pro was the most effective model using CWE-Sys2 + CWE-UserF setting. This exhibited the highest accuracy ($>$ 0.70) and $F_{1}$ (0.62).
In C, GPT-4 Turbo obtained balanced precision and recall (0.67 and 0.69), which resulted in the highest $F_{1}$ score among all six LLMs using the CWE-Sys2 + CWE-UserF prompt set. Moreover, similar to the result in RQ1, GPT-4o achieved a remarkable recall of 0.76.
For C++, GPT-4o consistently outperformed other models across precision, recall, and $F_{1}$ score, achieving the highest $F_{1}$ score (0.64) and recall (0.76) with both prompts for the few-shot setting.
Regarding JavaScript, GPT-4 Turbo was the most effective LLM using the CWE-Sys2 + CWE-UserF prompt configuration, especially in accuracy, precision, and $F_{1}$ score.
Lastly, GPT-4o demonstrated the best overall effectiveness, with the highest $F_{1}$ score (0.64) and precision (0.63) using CWE-Sys2 + CWE-UserF. Across all languages, the few-shot learning approach (CWE-UserF) consistently yielded better results compared to zero-shot prompts.
RQ2 (CWE classification): GPT-4o achieved the best vulnerability classification scores across multiple languages ($recall=0.74$ and $F_{1}=0.63$). GPT-4 Turbo showed strong performance in C and JavaScript. Gemini 1.5 Pro outperformed in Python. Few-shot prompts consistently outperformed zero-shot approaches across all models and languages.
IV CodeGuardian
To help engineers understand potential vulnerabilities in source code within their development workflow, we developed CodeGuardian, a VSCode extension that features just-in-time LLM-powered vulnerability scanning. It supports developers by providing real-time vulnerability analysis of their source code. More specifically, a user can highlight code snippets in editor, and CodeGuardian provides instant vulnerability analysis of the highlighted code. This feature helps developers identify and address security flaws at a snippet level. Multi-snippet or multi-file analyses are not supported.
Our study identified GPT-4 Turbo and GPT-4o as the best models for CWE detection and classification (see Section III). Consequently, CodeGuardian provides users with the option to select these models, using Azure OpenAI Studio APIs in the back end. The back-end system receives the highlighted code from a user and processes it to be concatenated to a specific prompt. After the vulnerability analysis, the results are shown on the extension panel.
The vulnerability analysis in CodeGuardian comprises two stages: first, a vulnerability detection phase, to determine whether the code is vulnerable or non-vulnerable, and second, a CWE classification phase, employing a few-shot setting for vulnerable code. Finally, the output of CodeGuardian is a detailed analysis of the CWE classification results.
V User Study
We evaluated CodeGuardian by a dedicated research question, with the main objective of providing a qualitative assessment of LLMs for security vulnerability detection.
RQ3 (usefulness): Does CodeGuardian improve developers’ accuracy and efficiency in vulnerability detection?
To address RQ3, we performed a user study within an industrial setting, which we describe next.
V-A Experiment Design
Participants and Grouping. Our experiments for the user study involved a diverse group of participants who have industrial software engineering experience. A total of 22 participants were selected, and we asked them to answer our pre-questionnaire. The pre-questionnaire contains several questions asking each participant about their experience with software development, software development skills, knowledge of cybersecurity, and familiarity with five selected programming languages. We used a 5-point Likert scale for the skills and knowledge questions and a nominal scale for the rest of the questions.
Participants were divided into two groups based on their responses: an experimental group (using CodeGuardian) and a control group (using only Internet search). Each group consists of 11 participants. We mainly referred to each participant’s years of experience in software development and cybersecurity to allocate them into equally distributed groups.
Challenges Design. All participants were required to perform a set of challenges during the experiment, which was identifying possible software vulnerabilities. In this experiment, a total of six challenges across five programming languages were presented to both groups. Table VI shows all the code snippets used for this user study, each snippet contains a specific vulnerability. While five snippets were vulnerable, one of the Python codes was used as a Placebo (i.e., not vulnerable).
Table VI: An overview of the six challenges of our user study.
No Language Vulnerability
1 C CWE-125: Out-of-bounds Read
2 C++ CWE-78: OS Command Injection
3 Java CWE-190: Integer Overflow or Wraparound
4 JavaScript CWE-22: Path Traversal
5 Python Placebo (Not Vulnerable)
6 Python CWE-79: Cross-site Scripting
Independent Variable. The independent variable was the type of tool used by the participants: CodeGuardian (LLM-based) and Internet search/own knowledge.
Dependent variables. The experiment for this user study has two dependent variables: (1) challenge completion time and (2) challenge completion accuracy.
Experimental Procedure. The first author of this study demonstrated how to use CodeGuardian exclusively for the experimental group, taking approximately 2-3 minutes to explain the usage instructions. Each participant then evaluated six code snippets, identifying any vulnerabilities and naming them or providing their CWE labels. The experimental group used CodeGuardian to address these challenges, while the control group could only use Internet resources, except for any LLM-based application. Participants had a maximum of 5 minutes to answer each question; if they finished early, the session progressed to the next task. All participants worked on a dedicated workstation to ensure consistent environmental conditions for all participants. This workstation was an EC2 instance running Debian 12 OS, accessible to all participants via a browser-based VNC.
Qualitative Analysis. The participants in the experimental group were asked to answer a post-questionnaire for the qualitative analysis of CodeGuardian. The questions focused on assessing the user experience and effectiveness of our plugin.
Data Analysis. To compare dependent variables across the experimental group and the control group, we analyzed the results using the Mann-Whitney U test [39] (with $\alpha=0.05$) and the magnitude of the differences using the rank-biserial correlation to calculate the effect size [40]. We selected this non-parametric test as completion time and accuracy data were not normally distributed.
V-B RQ3 (usefulness)
Figure 1: Challenge completion accuracy between the experimental group (green) and control group (yellow).
V-C Completion Accuracy
Accuracy was calculated based on the amount of correct answers out of all six challenges. Figure 1 visualizes the boxplot of challenge completion accuracy for both groups. We used a one-sided research hypothesis since we assumed that the experimental group had a higher percentage of correct answers than the control group. Therefore, the null and alternative hypotheses were defined as follows:
  • $H_{0}$: The two groups are the same in terms of the total challenge completion accuracy.
  • $H_{1}$: Total challenge completion accuracy in the experiment group was higher than in the control group.
The experimental group achieved a median accuracy of 1.0 (std = 0.08), while the control group had a median accuracy of 0.33 (std = 0.25). This represents a 203% increase in accuracy due to the use of CodeGuardian. Additionally, the smaller standard deviation in the experimental group’s accuracy indicates less variation in correct answer rates among its participants compared to the control group. This suggests that CodeGuardian not only enhanced overall accuracy but also contributed to more consistent performance.
The results of the Mann-Whitney U test indicated a statistically significant difference between the two groups (p-value=0.0001) with a large effect size (0.88). Thus, we reject the null hypothesis and support the alternative hypothesis, in which the experimental group achieved higher accuracy in completing the challenges. This indicates that the usage of CodeGuardian had a strong positive effect on improving the number of correct answers among participants.
V-D Completion Time
We measured the total challenge completion time (minutes:seconds) and compared the results statistically between the experimental and control groups. A one-sided research hypothesis was used for this test, as we hypothesized that the experimental group would complete all the challenges faster than the control group. Hence, the null and alternative hypotheses were defined as follows:
  • $H_{0}$: The two groups are the same in terms of the total challenge completion time.
  • $H_{1}$: Total challenge completion time in the experiment group is less than in the control group.
Figure 2 depicts the boxplots of completion time over all the challenges. The median and standard deviation in the experimental group were approximately 10:03 (mm:ss), and the standard deviation was 4:29 (mm:ss). On the contrary, the control group exhibited a mean of 24:59 (mm:ss) and a standard deviation of 4:58 (mm:ss). Although the standard deviation of both groups was almost identical, the experimental group completed all the challenges in 66% less time than the control group on average (around 15 minutes, on average).
The results of the Mann-Whitney U test showed a statistically significant difference (p-value = 0.00018) with a large effect size (0.95). Therefore, we reject the null hypothesis and conclude that the experimental group completed the challenges faster than the control group.
Figure 2: Challenge completion time between the experimental group (green) and control group (yellow).
Furthermore, Figure 3 illustrates the completion time per challenge. The median completion time in the experimental group was less than that of the control group in all challenges. Moreover, challenge 5 had the largest difference in performance (completion time) between the experimental and control groups by approximately 3.5 minutes. In contrast, challenge 2 in the experimental group showed a slight median difference compared to the control group($M_{exp}$=2:14 and $M_{ctl}$=3:20). Additionally, the lowest observed standard deviation for the experimental group was 46 seconds (Challenge 4), whereas the highest was 1 minute and 13 seconds (Challenge 1). Similar to the experimental group, the lowest observed median and standard deviation for the control group were 3:20 (Challenge 2) and 50 seconds (Challenge 6), and the highest median and standard deviation were 5 minutes (Challenge 1 and 6) and 1 minute and 34 seconds (Challenge 1). Although there were a few outliers in both experimental and control groups, the gaps between the lowest and highest standard deviations were relatively similar. Additionally, the experimental group completed the challenges on average by approximately 47% (Challenge 1), 44% (challenge 2), 58% (challenge 3), 59% (challenge 4), 65% (challenge 5), and 55% (challenge 6) faster, compared to the control group.
Figure 3: Comparison of challenge completion time per challenge between the experimental group (green) and control group (yellow).
RQ3 (usefulness): Both statistical tests for the challenge completion time and accuracy showed statistically significant improvements in the experimental group compared to the control group. CodeGuardian successfully increased the speed and accuracy of completing the challenges with large effect size.
VI Threats to Validity
VI-1 Internal validity
We compared all LLMs under identical experimental settings and on the same benchmark. Our implementation of the testing scripts to evaluate the scores was tested thoroughly. Moreover, we evaluated the LLMs with a limited set of prompts.
Our user study involved 22 participants, a reasonable number due to selection constraints (i.e., prior industrial knowledge, experience in software development, and cybersecurity). Although the sample size was small, it achieved the necessary statistical test power. The varying levels of software development and cybersecurity expertise among our participants, including their familiarity with programming languages and VSCode, could have influenced our results.
To mitigate this threat, we collected relevant background information through a pre-questionnaire and divided the participants into two groups, striving for an equal distribution of expertise. Also, we used the same environment for all participants (a workstation with Debian 12 OS and a connection speed of around 700 Mbps). However, the response time of the GPT-4 Turbo API was occasionally slower, especially during periods of high usage due to the limited number of regions where the model is hosted on Azure OpenAI Studio. The slower response times could potentially impact user experience and the seamless integration of CodeGuardian into the development workflow. These issues were reflected in the users’ responses to the post-questionnaire. However, it is important to note that this is a limitation of the current hosting infrastructure on Azure OpenAI Studio.
VI-2 External validity
The limited number of LLMs in our evaluation poses a threat in terms of the generalizability of our results, especially with the availability of newer models. Our results may not generalize or generalize differently when considering other LLMs. To mitigate this issue, we selected the best models available as of 05.2024. Likewise, our paper only involved five programming languages (C, C++, Python, Java, and JavaScript). This limited scope might not cover the diversity of programming environments and their associated vulnerabilities.
Furthermore, we used a relatively small number of code challenges in our user study. This was due to the limited availability of our participants, most of whom were industry professionals with constrained time. As a result, we included only one snippet per language, with an additional placebo snippet for Python. While this approach provided valuable insights into the effectiveness of CodeGuardian for vulnerability analysis, it may not fully represent all vulnerabilities and coding scenarios encountered in real-world projects.
VII Discussion
VII-A Model Selection
GPT-4 Turbo often exhibited higher effectiveness compared to the other large language models (see Section III). Although GPT-4o is the successor to GPT-4 Turbo, it did not consistently outperform its predecessor. GPT-4o achieved higher recall across all prompt setups, indicating it could identify more true positives. However, overall, GPT-4 Turbo excelled in three specific prompt setups (VD-Sys1 + VD-User2, VD-Sys2 + VD-User1, and VD-Sys2 + VD-User) primarily due to GPT-4o’s lower precision. This highlights a crucial aspect of using LLMs for vulnerability detection: newer versions or successors do not always outperform older models. Therefore, model selection should be based on specific security priorities. For instance, GPT-4o is preferable for detecting all potential vulnerabilities (i.e., reducing false negatives), whereas GPT-4 Turbo is better for minimizing false alarms. This underscores the importance of selecting the right model based on the particular requirements.
VII-B Model Performance in Each Language
Our study revealed significant variations in LLM performance across different programming languages for both vulnerability detection and CWE classification tasks. GPT-4-based models mostly outperformed other models across all languages. However, the strong performance of Gemini 1.5 Pro in JavaScript (vulnerability detection) and Python (CWE classification) indicates that different models may have strengths in particular languages or types of code analysis. This underscores the importance of considering language-specific performance when we select for secure code review.
VII-C Complexity of CWE Classification using LLM
Compared to vulnerability detection, our experiments reveal significant challenges in CWE classification, with performance varying across different CWE classes and model sizes. Larger models like GPT-4o, GPT-4 Turbo, and Gemini 1.5 Pro consistently outperformed smaller ones, particularly when employing few-shot learning techniques. For instance, we observed that recall rates of GPT-4o approximately triple using the few-shot setting. This improvement suggests that providing relevant examples helps models better understand the classification task. However, using few-shot learning does not help for smaller models, such as CodeLlama models due to the limited parameter size restricting their ability to effectively utilize the few-shot examples.
VII-D User Study
In our user study, the experimental group consistently completed challenges faster than the control group. Specifically, our placebo code analysis (Challenge 5) showed that most participants who used CodeGuardian were able to answer in under approximately two minutes. However, the control group’s participants took, on average, almost all of the allotted time (5 minutes). Thus, CodeGuardian helped participants quickly identify whether the code was vulnerable or not. We believe that this effect is because of the two-stage processing of CodeGuardian (see Section IV).
Furthermore, in Challenge 4, the most considerable reduction in completion time was found when participants used CodeGuardian. Analysis of our pre-questionnaire revealed that the participants were less skilled in JavaScript than in popular programming languages, such as Java or Python. Hence, we assume that, although some participants are not familiar with certain languages, CodeGuardian enables developers to analyze code and identify potential vulnerabilities regardless of prior knowledge.
In addition, accuracy was notably higher in the experimental group than in the control group. Evidence suggests that seamless, instant feedback and automated vulnerability analysis by CodeGuardian not only accelerate the analysis process but also correct vulnerability identification. Despite the majority of participants having limited knowledge of cybersecurity, according to the pre-questionnaire, the experimental group had an accuracy of more than 0.90 on average. In contrast, the accuracy of the control group was less than 0.50. This indicates that CodeGuardian effectively supports developers in identifying potential software vulnerabilities without having prior cybersecurity knowledge. However, our tool requires strict usage guidelines for junior developers, who might overly depend on their outputs without proper verification. Indeed, LLMs are prone to hallucinations, where the output appears plausible but is largely incorrect.
We believe CodeGuardian could have significant implications for secure software development practices. Reducing the time and effort required to identify vulnerabilities allows developers to conduct security checks more seamlessly within their workflow, potentially leading to more secure code being produced at a faster rate. The feedback from the experimental group indicated that the majority of participants ($>$ 70%) were satisfied with the speed of the vulnerability scanning process. Furthermore, users particularly appreciated how CodeGuardian effectively checked potential vulnerabilities automatically.
VIII Related Work
VIII-A Conventional Methods and Tools
Conventional vulnerability analysis primarily relies on static application security testing (SAST) and dynamic application security testing (DAST). SAST examines source code without execution and focuses on syntactic and semantic checks. DAST evaluates applications during runtime, mimicking attacker behavior [41]. However, SAST typically generates a high number of false positives [42], and DAST provides high false negatives, which may miss many vulnerabilities [43].
In constrast, in this paper we focus on LLM-based approaches to vulnerability detection.
VIII-B ML and DL in Vulnerability Detection
Recently, vulnerability detection systems have become more adapted to learning from data to predict and identify vulnerabilities with higher accuracy [16]. DL approaches, particularly utilizing Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks, have further improved detection accuracy without requiring feature engineering by hand [15, 44]. Despite the advancement brought by DL, these models treat code as linear sequences, potentially overseeing complex semantic relationships as well as long-range dependencies [45].
These approaches are challenged by the development of pre-trained LLMs. Unlike DL approaches, LLMs such as the ones used in this paper require no explicit training, while achieving high scores in vulnerability detection.
VIII-C LLMs in Vulnerability Detection and Classification
Studies have shown that fine-tuned models like BERT and GPT outperform traditional DL models in vulnerability detection [10, 9]. However, this approach requires substantial computational resources and large, well-annotated datasets. In contrast, prompt engineering leverages pre-trained LLMs without any parameter update, offering a more resource-efficient approach. Several studies have demonstrated the effectiveness of prompt engineering, such as few-shot learning and chain-of-thought in vulnerability detection and classification [46, 47, 48]. Specifically, Zhou et al. [37] showed that GPT-4, using prompt engineering, outperformed fine-tuned CodeBERT.
Although prompt engineering does not require a large dataset, the effectiveness of LLM in vulnerability detection and classification heavily relies on the quality of the dataset used for evaluation. Recent studies highlighted data quality issues in existing datasets, such as poor data quality, low label accuracy, and high data duplication rates [26, 27], causing misleading results when measuring the effectiveness of LLMs.
Since this study focused on optimizing input prompts and inference methods without fine-tuning, it was important to ensure the use of a high-quality dataset. Moreover, to the best of our knowledge, existing works are limited in terms of language diversity, which is a characterizing element of our work.
VIII-D Tools Leveraging LLMs for Vulnerability Analysis
Although ChatGPT is not specifically designed to analyze and detect vulnerabilities, it has shown promising results [49, 14]. However, it is challenging to integrate ChatGPT seamlessly into the development workflow. To address this limitation, researchers have also developed IDE-integrated solutions [50, 51]. Yet, current tools rely on less advanced LLMs, potentially limiting their effectiveness compared to state-of-the-art LLMs like GPT-4.
To this aim, we developed CodeGuardian as a widely accessible VSCode extension that makes state of the art LLM models accessible. Our user study with industrial participants suggested the benefits and perils of a more pervasive adoption of LLMs for vulnerability detection beyond purely academic studies, a novel contribution to the state of the art.
IX Conclusions and Future Work
This paper presented a comprehensive study on the effectiveness of large language models (LLMs) for secure code assessment across multiple programming languages. We evaluated six state-of-the-art LLMs on vulnerability detection and CWE classification tasks using our developed multi-language vulnerability dataset. Our results demonstrate that GPT-4 Turbo and GPT-4o outperform other LLMs. Although the GPT-4-based models often achieved the highest effectiveness in both tasks and different programming languages, Gemini 1.5 Pro showed strong performance in JavaScript and Python.
Furthermore, we found that few-shot learning techniques significantly improve performance, especially for CWE classification. To bridge the gap between research findings and practical applications, we developed CodeGuardian integrating LLM-powered vulnerability analysis into developers’ workflows. Our user study with 22 industry-experienced developers showed that CodeGuardian significantly improves both the speed and accuracy of vulnerability detection, with participants completing tasks 66% faster and achieving 203% higher accuracy compared to traditional methods.
For future work, we should focus on expanding our manually annotated dataset to improve robustness and generalizability. Exploring advanced prompt engineering methods such as chain-of-thought and Retrieval Augmented Generation (RAG) could further enhance LLM performance. Additionally, improving CodeGuardian’s user interface and incorporating user feedback will make it a more effective tool for secure software development.

### v2

@@
 Large Language Models for Secure Code Assessment: A Multi-Language Empirical Study
-Abstract
-Most vulnerability detection studies focus on datasets of vulnerabilities in C/C++ code, offering limited language diversity.
-Thus, the effectiveness of deep learning methods, including large language models (LLMs), in detecting software vulnerabilities beyond these languages is still largely unexplored.
-In this paper, we evaluate the effectiveness of LLMs in detecting and classifying Common Weakness Enumerations (CWE) using different prompt and role strategies.
-Our experimental study targets six state-of-the-art pre-trained LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLLama-7B, CodeLLama-13B, and Gemini 1.5 Pro) and five programming languages: Python, C, C++, Java, and JavaScript.
-We compiled a multi-language vulnerability dataset from different sources, to ensure representativeness.
-Our results showed that GPT-4o achieves the highest vulnerability detection and CWE classification scores using a few-shot setting.
-Aside from the quantitative results of our study, we developed a library called CodeGuardian integrated with VSCode which enables developers to perform LLM-assisted real-time vulnerability analysis in real-world security scenarios.
-We have evaluated CodeGuardian with a user study involving 22 developers from the industry.
-Our study showed that, by using CodeGuardian, developers are more accurate and faster at detecting vulnerabilities.
-I Introduction
-A software vulnerability is a defect that could allow an attacker to gain control of a software system, steal or manipulate sensitive data, install a backdoor, or plant other types of malware [1, 2, 3].
-Software vulnerabilities vary in nature and level of exploitability.
-Not all vulnerabilities allow attackers to cause the same types of harm, and not all vulnerabilities create equally severe risks.
-However, all vulnerabilities pose at least some level of risk to the applications they impact, as well as the environments that host those applications.
-Software vulnerabilities may have different root causes, spanning from the way the software is designed, problems with the software’s source code, poor management of data, or access control settings within the application.
-In this work, we focus on classes of vulnerabilities related to bugs in the source code.
-The U.S.
-Department of Homeland Security reports that 90% of security incidents are caused by exploiting security flaws and defects in the code [4].
-A recent survey [5] found that over 55% of developers struggle to identify code vulnerabilities, indicating a lack of secure coding knowledge.
-Due to the increasing frequency and severity of cybersecurity incidents, which negatively impact end-users, businesses, and critical infrastructure [6], identifying security vulnerabilities within source code has become essential.
-Researchers have proposed methods for the automatic detection and repair of software vulnerabilities [7, 8, 9, 10, 11, 12, 13].
-For instance, program analysis-based techniques have been widely employed [14].
-However, such rule-based techniques are affected by high false positive rates and low effectiveness in handling a large number of diverse types of vulnerabilities [12].
-In addition, they often suffer from long response times [13], which makes them less suitable for continuous integration/development environments.
-Recently, Deep Learning (DL) techniques have enabled significant progress in vulnerability detection in the source code [15, 16, 17].
-However, most DL models treat the code as linear sequences or textual data, which limits the finding of the semantic relationship within the source code and demands the development of more robust and efficient methods.
-To address these limitations, researchers have started exploring the potential of Large Language Models (LLMs) that are pre-trained on a large corpus of data and use advanced natural language processing capabilities based on the transformer architecture [18].
-LLMs have shown great results across various natural language and software engineering tasks [18, 11], including automated vulnerability detection and repair [9, 10, 11].
-Existing studies on vulnerability detection are executed against datasets that include vulnerabilities pertaining to only one or two programming languages.
-Instances are the vulnerability-fixing commit dataset by Zhou et al.
-[19] (C/C++), by Li et al.
-[20] (Java), BigVul [21] (C/C++), SVEN [22] (Python and C/C++), Devign [23] (C), D2A [24] (C/C++).
-Consequently, the capabilities of LLMs across diverse programming languages remain largely unexplored.
-Furthermore, these studies are purely quantitative, and the potential of LLMs to assist developers in identifying vulnerabilities has not yet been investigated.
-To fill these research gaps, this paper investigates the effectiveness of LLMs in identifying vulnerable codes across different programming languages.
-To the best of our knowledge, this is the first study that reports an extensive evaluation of six state-of-the-art pre-trained LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLLama-7B, CodeLLama-13B, and Gemini 1.5 Pro) in identifying the vulnerabilities pertaining to five programming languages, namely Python, C, C++, Java, and JavaScript.
-Our study focuses on the top 25 Common Weakness Enumeration (CWE) classes that represent the most dangerous software weaknesses [25], for which we compiled a dataset of vulnerable snippets from existing reference datasets.
-Our study targets vulnerability detection and CWE classification using zero-shot and few-shot approaches.
-Furthermore, to investigate the industrial applicability of LLMs in software vulnerability detection and debugging, we have developed a VSCode extension called CodeGuardian.
-It integrates the LLM analyzed by our quantitative study to allow developers to scan code dynamically while editing, providing instant feedback and enhancing secure coding practices directly within their workflow.
-We have evaluated CodeGuardian with a user study involving six tasks across the five programming languages and 22 participants with software engineering experience in the industry and varying levels of expertise in software security.
-Our qualitative analysis revealed that by using CodeGuardian, developers were twice as accurate at CWE classification and 60% faster, thanks to the LLM-powered engine.
-Our paper makes the following contributions:
-Dataset.
-A dataset of more than 370 manually validated vulnerabilities, pertaining to five programming languages, which can support the evaluation of automated techniques in security vulnerability detection and classification.
-Evaluation. An empirical study concerning the capabilities of six pre-trained LLMs at detecting and classifying software vulnerabilities in five programming languages.
-CodeGuardian. A VSCode extension that allows developers to use LLMs to scan code for vulnerabilities while editing.
-User Study. A qualitative analysis of CodeGuardian with a user study of 22 developers from our industrial partner.
-II Empirical Study
-II-A Research Questions
+Abstract.
+Most vulnerability detection studies rely on datasets tied to specific programming languages and within-project settings, limiting language diversity and the assessment of model generalizability.
+Consequently, the effectiveness of deep learning methods, including large language models (LLMs), beyond such restricted scenarios remains largely unexplored.
+In this paper, we evaluate six state-of-the-art LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLlama-7B, CodeLlama-13B, Gemini 1.5 Pro) on vulnerability detection and CWE classification across five languages (Python, C, C++, Java, JavaScript).
+We compiled a multi-language dataset from diverse sources and investigated different prompt and role strategies.
+Our results show that GPT-4o achieves the best performance, particularly in few-shot settings.
+To examine practical applicability, we developed CodeGuardian, a VSCode extension that enables real-time LLM-assisted vulnerability detection.
+In a user study with 22 professional developers, CodeGuardian doubled accuracy and halved task completion time compared to a control group.
+Vulnerability Detection, Large Language Models, Security, LLMs
+journalyear: 2026copyright: ccconference: 7th International Workshop on Deep Learning for Testing and Testing for Deep Learning (DeepTest ’26); April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 7th International Workshop on Deep Learning for Testing and Testing for Deep Learning (DeepTest ’26) (DeepTest ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786154.3788584isbn: 979-8-4007-2386-5/2026/04ccs: Software and its engineering Software verification and validation
+[…]
+by
+1. Introduction
+Software vulnerabilities are defects that can enable attackers to steal data, gain control of systems, or insert malware (Shirey, 2007; Shoshitaishvili et al., 2015; Or-Meir et al., 2019).
+They arise from design flaws, insecure code, or poor data and access management.
+In this work, we focus on vulnerabilities rooted in source code bugs.
+According to the U.S.
+Department of Homeland Security, 90% of security incidents exploit such flaws (Cybersecurity and Infrastructure Security Agency (CISA), 2021).
+Yet, over half of developers report difficulties identifying vulnerabilities (GitLab, 2022), highlighting the need for effective detection tools.
+Traditional program-analysis techniques (Pistoia et al., 2007; Pinconschi et al., 2021; Thapa et al., 2022; Alqarni and Azim, 2022; Chen et al., 2023b; Johnson et al., 2013; Rangnau et al., 2020) are limited by high false positives, narrow coverage, and long analysis times (Johnson et al., 2013; Rangnau et al., 2020).
+Deep learning approaches (Russell et al., 2018; Hanif et al., 2021; Dam et al., 2021; Humbatova et al., 2020) improved scalability but often treat code as text, failing to capture semantic structure (Humbatova et al., 2020; Riccio et al., 2020).
+Large Language Models (LLMs), pre-trained on massive code and text corpora (Vaswani et al., 2017), have recently shown promise for vulnerability detection (Thapa et al., 2022; Alqarni and Azim, 2022; Chen et al., 2023b).
+However, existing evaluations are limited to one or two languages (e.g., BigVul (Fan et al., 2020), SVEN (He and Vechev, 2023)), and LLM capabilities across diverse languages remain largely unexplored.
+Moreover, most studies remain purely quantitative, with little evidence on how LLMs can assist developers in practice.
+This paper addresses these gaps by evaluating six state-of-the-art LLMs (GPT-3.5-Turbo, GPT-4 Turbo, GPT-4o, CodeLlama-7B, CodeLlama-13B, Gemini 1.5 Pro) on vulnerability detection and CWE classification across five languages (Python, C, C++, Java, JavaScript).
+We focus on the top 25 CWE classes (1) and compile a dataset of over 370 manually validated vulnerable snippets.
+Beyond our quantitative analysis, we investigate practical applicability through CodeGuardian, a VSCode extension that integrates LLMs into developers’ workflows.
+In a user study with 22 industry practitioners, CodeGuardian doubled accuracy in CWE classification and reduced task completion time by 60%.
+The contributions of this work are as follows:
+Dataset. A curated dataset of 370+ vulnerabilities across five languages, supporting evaluation of detection and classification tools.
+Evaluation. An empirical study of six LLMs on vulnerability detection and classification across multiple languages.
+CodeGuardian. A VSCode extension enabling real-time vulnerability scanning, released in our replication package (rep-package, 2025).
+User Study. A qualitative evaluation with 22 professional developers, showing significant improvements in speed and accuracy with CodeGuardian.
+2. Empirical Study
+2.1. Research Questions
 Our study considers the following research questions:
 RQ1 (vulnerability detection): How effective are LLMs in detecting CWEs across multiple languages? How does effectiveness vary among different languages?
-RQ2 (CWE classification): How effective are LLMs in classifying CWEs across multiple languages? How does effective vary across languages? How does effectiveness vary among different languages?
-The first research question evaluates how accurately LLMs perform binary classification, determining whether a code snippet is vulnerable or not.
-The second research question evaluates the accuracy of LLMs in classifying CWEs by assigning the correct CWE ID to a given code snippet.
-II-B Dataset Design
-The representativeness of datasets is crucial for effective vulnerability detection and classification [26, 27].
-Existing works adopted datasets with limited language diversity, focusing primarily on C/C++ [19, 20, 21, 22, 23, 24].
-In contrast, in this work, we consider software vulnerabilities of five popular programming languages, namely Python, C, C++, Java, and JavaScript.
-We compiled a dataset that integrates three existing vulnerability datasets: CVEFixes, CWE-snippets, and the JavaScript Vulnerability DataSet (JVD).
-Table I shows the statistics of these datasets, including the number of languages, CWEs, and files.
-CVEFixes is a fix-commit vulnerability dataset [28].
+RQ2 (CWE classification): How effective are LLMs in classifying CWEs across multiple languages? How does effectiveness vary among different languages?
+2.2. Dataset Design
+The representativeness of datasets is crucial for effective vulnerability detection and classification (Ding et al., 2024; Croft et al., 2023).
+Most existing works adopted datasets with limited language diversity, focusing primarily on C/C++ (Zhou et al., 2024a; Li et al., 2023; Fan et al., 2020; He and Vechev, 2023; Zhou et al., 2019; Zheng et al., 2021).
+Although some researchers have attempted to address language diversity limitations with evaluating LLMs on three to seven programming languages (Shu et al., 2025; Atiiq et al., 2024; Zhang et al., 2025), they either focus solely on binary vulnerability detection without CWE classification (Shu et al., 2025; Atiiq et al., 2024).
+Additionally, most existing studies rely heavily on publicly available vulnerability data primarily from the National Vulnerability Database (NVD), potentially limiting the diversity and real-world applicability of their findings.
+In contrast, in this work, we consider software vulnerabilities of five popular programming languages, namely Python, C, C++, Java, and JavaScript, which represent the top-ranked languages according to IEEE’s 2025 survey (IEEE Spectrum, 2025).
+Furthermore, we compiled a dataset that integrates three existing vulnerability datasets: CVEFixes, CWE-snippets, and the JavaScript Vulnerability DataSet (JVD).
+Table 1 shows the statistics of these datasets, including the number of languages, CWEs, and files.
+CVEFixes is a fix-commit vulnerability dataset (Bhandari et al., 2021).
 CWE-snippets is a dataset created in collaboration with an industrial partner for this study.
 It consists of numerous vulnerability snippets across multiple programming languages.
-JVD is a vulnerability dataset for JavaScript code [29].
-From each dataset, we selected the top 25 CWE most dangerous software weaknesses [25] for each of the five selected programming languages.
-Table I: Vulnerabilities benchmarks evaluated in this study.
+JVD is a vulnerability dataset for JavaScript code (Ferenc et al., 2019).
+From each dataset, we selected the top 25 CWE most dangerous weaknesses (1) for each of the five selected programming languages.
+Table 1. Vulnerabilities benchmarks evaluated in this study.
 Dataset # Languages # CWEs # Files
-      CVEFixes [28]       27       180       18,249
+CVEFixes (Bhandari et al., 2021) 27 180 18,249
 CWE-snippets 42 506 231,571
-      JVD [29]       1       N/A       12,125
-The maximum number of snippets per class was set to three across all the languages due to a limited number of vulnerable code samples available in the reference datasets for certain CWE classes and programming languages.
-For example, CWE-787 (Out-of-bounds Write) is more commonly associated with memory-unsafe languages like C and C++, while it is less prevalent in memory-safe languages such as Python and Java.
-Consequently, we prioritized constructing a balanced representation of vulnerabilities across all languages in our dataset.
-As a result, we aimed to construct a balanced representation of vulnerabilities across all languages in our dataset.
-In total, we retained 378 snippets, ensuring an equal number of vulnerable and non-vulnerable snippets for each programming language ($n_{Python}=38$, $n_{Java}=42$, $n_{C++}=36$, $n_{C}=44$, and $n_{JavaScript}=29$).
-Table II: LLMs evaluated for vulnerability detection and classification.
+JVD (Ferenc et al., 2019) 1 N/A 12,125
+2.2.1. Dataset Construction and Validation
+Given the critical importance of data quality in vulnerability detection research (Ding et al., 2024), we adopted a systematic approach for dataset construction that addresses known limitations in existing vulnerability datasets.
+Our methodology involves several key phases designed to ensure both representativeness and accuracy of the collected data.
+The data collection process involved executing systematic queries against the CVEFixes and CWE-snippets databases to extract code snippets corresponding to specific CWE categories across all target programming languages.
+Since JVD was structured as a CSV file, we employed systematic filtering techniques to identify relevant JavaScript vulnerabilities.
+To maintain dataset balance and address the inherent scarcity of certain vulnerability types in specific languages, we set a maximum threshold of three snippets per CWE class across all languages.
+This constraint was particularly important given that certain vulnerabilities—such as CWE-787 (Out-of-bounds Write)—tend to be associated with memory-unsafe languages like C and C++, while remaining relatively uncommon in memory-safe languages such as Python and Java.
+Furthermore, our methodology involved manual validation of all selected code snippets.
+Each snippet has been inspected to verify both the presence of the vulnerability and the associated CWE ID.
+This validation process was essential to mitigate the data quality issues that have been identified in existing vulnerability datasets (Ding et al., 2024).
+For binary vulnerability detection tasks, we curated non-vulnerable code snippets from multiple sources within our reference datasets.
+These included fixed versions of vulnerable code from CVEFixes commit histories, solution implementations from the CWE-snippets dataset, and verified secure code samples from JVD.
+The selection process for non-vulnerable snippets required careful manual inspection to ensure they represented genuinely secure implementations rather than simply alternative vulnerable patterns.
+It is important to note that the majority of non-vulnerable snippets in our dataset are not direct fixed versions of the corresponding vulnerable snippets.
+Based on our manual inspection, only a limited number of snippet pairs existed where the non-vulnerable code was the direct fixed version of a specific vulnerable snippet (one pair in C++ and two pairs in C).
+The use of independent non-vulnerable samples would also support model generalizability by ensuring exposure to diverse secure coding patterns rather than specific fix implementations.
+Our final dataset contains 378 manually validated code snippets, balanced across vulnerable and non-vulnerable samples for each language (Table 2).
+It includes Python (38/38), Java (42/42), C++ (36/36), C (44/44), and JavaScript (29/29).
+Most samples were drawn from the CWE-snippets and CVEFixes datasets, while JVD provided the non-vulnerable JavaScript cases.
+To account for the hierarchical nature of CWE, predictions matching a child CWE (e.g., CWE-78) for a snippet labeled with its parent (e.g., CWE-77) were considered correct, with the parent assigned as the final class.
+Table 2. Number of code snippets (vul. / non-vul.).
+Language CVE fixes CWE snippets JSV Total
+C 15 / 7 29 / 37 0 / 0 44 / 44
+C++ 12 / 15 24 / 21 0 / 0 36 / 36
+Java 16 / 3 26 / 39 0 / 0 42 / 42
+JavaScript 15 / 4 14 / 2 0 / 23 29 / 29
+Python 22 / 38 16 / 0 0 / 0 38 / 38
+Total 80 / 67 109 / 99 0 / 23 189 / 189
+Table 3. LLMs evaluated in this work.
 Model Parameters Architecture Input Tokens Output Tokens
-GPT-3.5-Turbo [30] $>$ 100B Decoder 16,385 4,096
-GPT-4 Turbo [31] $>$ 100B Decoder 128K 4,096
-GPT-4o [32] $>$ 100B Decoder 128K 4,096
-CodeLLama-7B [33] 7 B Decoder 100K 4,096
-CodeLLama-13B [33] 13 B Decoder 100K 4,096
-Gemini 1.5 Pro [34] $>$ 100B Decoder 1M 8,192
-Table III: System and user prompts: VD-Sys1, VD-Sys2, VD-User1 and VD-User2 are used for vulnerability detection (RQ1).
+GPT-3.5-Turbo (OpenAI, 2024a) >> 100B Decoder 16,385 4,096
+GPT-4 Turbo (OpenAI, 2024b) >> 100B Decoder 128K 4,096
+GPT-4o (OpenAI, 2024c) >> 100B Decoder 128K 4,096
+CodeLLama-7B (rozière2024codellamaopenfoundation) 7B Decoder 100K 4,096
+CodeLLama-13B (rozière2024codellamaopenfoundation) 13B Decoder 100K 4,096
+Gemini 1.5 Pro (Team, 2024) >> 100B Decoder 1M 8,192
+Table 4.
+System and user prompts: Sys1, VD-Sys2, VD-User1 and VD-User2 are used for vulnerability detection (RQ1).
 CWE-Sys1 and CWE-Sys2 are used for system prompts 1 and 2 for CWE Classification.
@@
 No. Prompt Type Prompt Template Verbalizer
-VD-Sys1 System Description You are an AI binary vulnerability classifier that identifies whether the provided code is vulnerable or not vulnerable.
+V-Sys1 System Description You are an AI binary vulnerability classifier that identifies whether the provided code is vulnerable or not vulnerable.
 You should respond with either only ’vulnerable’ or ’not vulnerable’.
 N/A
-VD-Sys2 System Description You are an experienced developer who knows the security vulnerability very well. N/A
-VD-User1 User Description for VD Classify the following code in vulnerable or not vulnerable. Output either only ’vulnerable’ or ’not vulnerable’. +: vulnerable
--: non-vulnerable
-VD-User2 User Description for VD Now you need to identify whether a code contains a vulnerability or not. If has any potential vulnerability, output: ’vulnerable’. Otherwise, output: ’not vulnerable’.
+V-Sys2 System Description You are an experienced developer who knows the security vulnerability very well. N/A
+U1 User Description for VD Classify the following code in vulnerable or not vulnerable. Output either only ’vulnerable’ or ’not vulnerable’. vulnerable
+or
+non-vulnerable
+U2 User Description for VD Now you need to identify whether a code contains a vulnerability or not. If has any potential vulnerability, output: ’vulnerable’. Otherwise, output: ’not vulnerable’.
 You must respond with either ’vulnerable’ or ’not vulnerable’ only. The code is below.
-+: vulnerable
--: non-vulnerable
-CWE-Sys1 System Description for CWE You are an AI vulnerability classifier that identifies CWE of a sent code, if that code has a vulnerability.
+vulnerable
+or
+non-vulnerable
+C-Sys1 System Description for CWE You are an AI vulnerability classifier that identifies CWE of a sent code, if that code has a vulnerability.
 You should respond with only one CWE. You don’t have to respond with any other information except for one of the CWEs.
 N/A
-CWE-Sys2 System Description for CWE Classification You are an experienced developer who knows the security vulnerability very well. N/A
-CWE-UserZ User Description for CWE (zero-shot) Classify the following code in CWE a category.
+C-Sys2 System Description for CWE Classification You are an experienced developer who knows the security vulnerability very well. N/A
+UZ User Description for CWE (zero-shot) Classify the following code in CWE a category.
 Output only one CWE tag in lowercase letters.
@@
 CWE ID
-CWE-UserF User Description for CWE Classification with the few-shot setting Classify the following code in a CWE category.
+UF User Description for CWE Classification with the few-shot setting Classify the following code in a CWE category.
 The code should have a vulnerability that corresponds to one of the top-25 CWEs below.
@@
 CWE ID
-II-C Large Language Models
+2.3. Large Language Models
 We evaluate the effectiveness of using six state-of-the-art pre-trained LLMs.
-Table II shows the selected LLMs, including their parameter size (the number of parameters for GPT and Gemini 1.5 Pro models are not disclosed by the organizations, and therefore, we report an estimate), architecture, input, and output token sizes.
-Specifically, we considered GPT-3.5 Turbo, GPT-4 Turbo, and CodeLlama models available as of June 2024, and pertaining from different organizations such as OpenAI’s GPT-3.5-Turbo [30], GPT-4 Turbo [31] and GPT-4o [32], as well as Meta’s CodeLLama-7B [33] and CodeLLama-13B [33], and Google’s Gemini 1.5 Pro [34].
+Table 3 shows the selected LLMs, including their parameter size (the number of parameters for GPT and Gemini 1.5 Pro models are not disclosed by the organizations, and therefore, we report an estimate), architecture, input, and output token sizes.
+Specifically, we considered GPT-3.5 Turbo, GPT-4 Turbo, and CodeLlama models available as of June 2024, and pertaining from different organizations such as OpenAI’s GPT-3.5-Turbo (OpenAI, 2024a), GPT-4 Turbo (OpenAI, 2024b) and GPT-4o (OpenAI, 2024c), as well as Meta’s CodeLLama-7B (rozière2024codellamaopenfoundation) and CodeLLama-13B (rozière2024codellamaopenfoundation), and Google’s Gemini 1.5 Pro (Team, 2024).
 For the GPT models, we utilized the supporting company’s Azure OpenAI Studio.
-Concerning hosting CodeLLama-7B and CodeLLama-13B models, we employed Google Colab and the transformers library by HuggingFace [35].
+Concerning CodeLLama-7B and CodeLLama-13B models, we employed Google Colab and the transformers library by HuggingFace (et al., 2020).
 For the Gemini 1.5 Pro model, we used the Vertex AI from the Google Cloud Platform.
-Lastly, to minimize the models’ creativity and obtain more focused, conservative, and consistent responses, we set the temperature parameter to $0.1$ for all LLMs, as done in previous studies [36].
-Table IV: RQ1: Effectiveness comparison of six LLMs using four prompt configurations for CWE detection. The best scores are indicated in bold.
-LLM / Prompt VD-Sys1 + VD-User1 VD-Sys1 + VD-User2 VD-Sys2 + VD-User1 VD-Sys2 + VD-User2
-Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$
+Lastly, to minimize the models’ creativity and obtain more focused, conservative, and consistent responses, we set the temperature parameter to $0.1$ for all LLMs, as done in previous studies (Chen et al., 2023a).
+Table 5. RQ1: Effectiveness comparison of six LLMs using four prompt configurations for CWE detection.
+LLM / Prompt V-Sys1–U1 V-Sys1–U2 C-Sys2–U1 C-Sys2–U2
+Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
 All
-GPT-3.5 Turbo 0.62 0.59 0.77 0.67 0.57 0.55 0.82 0.66 0.72 0.73 0.70 0.71 0.61 0.58 0.77 0.66
-GPT-4 Turbo 0.81 0.81 0.82 0.81 0.79 0.73 0.92 0.81 0.79 0.75 0.87 0.81 0.78 0.71 0.93 0.81
-GPT-4o 0.80 0.72 0.97 0.83 0.71 0.64 0.99 0.78 0.62 0.57 1.00 0.72 0.63 0.57 0.99 0.73
-CodeLlama-7b 0.62 0.58 0.89 0.70 0.57 0.54 0.93 0.68 0.54 0.52 0.98 0.68 0.57 0.55 0.84 0.66
-CodeLlama-13b 0.60 0.76 0.28 0.41 0.66 0.67 0.65 0.66 0.65 0.62 0.77 0.69 0.66 0.62 0.81 0.71
-Gemini 1.5 Pro 0.80 0.79 0.83 0.81 0.69 0.65 0.84 0.73 0.75 0.71 0.83 0.77 0.68 0.62 0.97 0.75
+GPT-3.5 Turbo 0.59 0.77 0.67 0.55 0.82 0.66 0.73 0.70 0.71 0.58 0.77 0.66
+GPT-4 Turbo 0.81 0.82 0.81 0.73 0.92 0.81 0.75 0.87 0.81 0.71 0.93 0.81
+GPT-4o 0.72 0.97 0.83 0.64 0.99 0.78 0.57 1.00 0.72 0.57 0.99 0.73
+CodeLlama-7b 0.58 0.89 0.70 0.54 0.93 0.68 0.52 0.98 0.68 0.55 0.84 0.66
+CodeLlama-13b 0.76 0.28 0.41 0.67 0.65 0.66 0.62 0.77 0.69 0.62 0.81 0.71
+Gemini 1.5 Pro 0.79 0.83 0.81 0.65 0.84 0.73 0.71 0.83 0.77 0.62 0.97 0.75
 Python
-GPT-3.5 Turbo 0.63 0.59 0.87 0.70 0.57 0.54 0.89 0.67 0.74 0.75 0.71 0.73 0.61 0.57 0.87 0.69
-GPT-4 Turbo 0.75 0.73 0.79 0.76 0.78 0.72 0.89 0.80 0.78 0.73 0.87 0.80 0.71 0.66 0.87 0.75
-GPT-4o 0.76 0.70 0.92 0.80 0.66 0.60 0.95 0.73 0.58 0.54 1.00 0.70 0.59 0.55 0.97 0.70
-CodeLlama-7b 0.70 0.67 0.76 0.72 0.58 0.56 0.79 0.65 0.54 0.52 0.97 0.68 0.57 0.55 0.76 0.64
-CodeLlama-13b 0.57 0.69 0.24 0.35 0.59 0.59 0.61 0.60 0.59 0.59 0.63 0.61 0.54 0.53 0.63 0.58
-Gemini 1.5 Pro 0.76 0.79 0.71 0.75 0.61 0.59 0.71 0.64 0.74 0.70 0.84 0.76 0.66 0.60 0.95 0.73
+GPT-3.5 Turbo 0.59 0.87 0.70 0.54 0.89 0.67 0.75 0.71 0.73 0.57 0.87 0.69
+GPT-4 Turbo 0.73 0.79 0.76 0.72 0.89 0.80 0.73 0.87 0.80 0.66 0.87 0.75
+GPT-4o 0.70 0.92 0.80 0.60 0.95 0.73 0.54 1.00 0.70 0.55 0.97 0.70
+CodeLlama-7b 0.67 0.76 0.72 0.56 0.79 0.65 0.52 0.97 0.68 0.55 0.76 0.64
+CodeLlama-13b 0.69 0.24 0.35 0.59 0.61 0.60 0.59 0.63 0.61 0.53 0.63 0.58
+Gemini 1.5 Pro 0.79 0.71 0.75 0.59 0.71 0.64 0.70 0.84 0.76 0.60 0.95 0.73
 C
-GPT-3.5 Turbo 0.56 0.55 0.68 0.61 0.47 0.48 0.70 0.57 0.69 0.68 0.73 0.70 0.57 0.56 0.68 0.61
-GPT-4 Turbo 0.83 0.78 0.91 0.84 0.75 0.67 0.98 0.80 0.80 0.75 0.89 0.81 0.78 0.70 0.98 0.82
-GPT-4o 0.83 0.75 1.00 0.85 0.70 0.63 1.00 0.77 0.64 0.58 1.00 0.73 0.66 0.59 1.00 0.75
-CodeLlama-7b 0.59 0.55 0.93 0.69 0.59 0.55 0.98 0.70 0.52 0.51 0.95 0.67 0.52 0.51 0.86 0.64
-CodeLlama-13b 0.65 0.93 0.32 0.47 0.69 0.67 0.77 0.72 0.65 0.60 0.86 0.71 0.77 0.72 0.89 0.80
-Gemini 1.5 Pro 0.77 0.73 0.86 0.79 0.70 0.67 0.80 0.73 0.78 0.79 0.77 0.78 0.73 0.65 1.00 0.79
+GPT-3.5 Turbo 0.55 0.68 0.61 0.48 0.70 0.57 0.68 0.73 0.70 0.56 0.68 0.61
+GPT-4 Turbo 0.78 0.91 0.84 0.67 0.98 0.80 0.75 0.89 0.81 0.70 0.98 0.82
+GPT-4o 0.75 1.00 0.85 0.63 1.00 0.77 0.58 1.00 0.73 0.59 1.00 0.75
+CodeLlama-7b 0.55 0.93 0.69 0.55 0.98 0.70 0.51 0.95 0.67 0.51 0.86 0.64
+CodeLlama-13b 0.93 0.32 0.47 0.67 0.77 0.72 0.60 0.86 0.71 0.72 0.89 0.80
+Gemini 1.5 Pro 0.73 0.86 0.79 0.67 0.80 0.73 0.79 0.77 0.78 0.65 1.00 0.79
 C++
-GPT-3.5 Turbo 0.60 0.57 0.75 0.65 0.60 0.56 0.86 0.68 0.72 0.71 0.75 0.73 0.61 0.59 0.75 0.66
-GPT-4 Turbo 0.85 0.82 0.89 0.85 0.83 0.76 0.97 0.85 0.78 0.71 0.94 0.81 0.81 0.73 0.97 0.83
-GPT-4o 0.86 0.78 1.00 0.88 0.75 0.67 1.00 0.80 0.62 0.57 1.00 0.73 0.60 0.55 1.00 0.71
-CodeLlama-7b 0.61 0.57 0.92 0.70 0.57 0.54 0.94 0.69 0.57 0.54 1.00 0.70 0.60 0.56 0.94 0.70
-CodeLlama-13b 0.61 0.83 0.28 0.42 0.62 0.63 0.61 0.62 0.68 0.64 0.83 0.72 0.69 0.64 0.89 0.74
-Gemini 1.5 Pro 0.88 0.85 0.92 0.88 0.72 0.67 0.89 0.76 0.72 0.67 0.89 0.76 0.68 0.61 0.97 0.75
+GPT-3.5 Turbo 0.57 0.75 0.65 0.56 0.86 0.68 0.71 0.75 0.73 0.59 0.75 0.66
+GPT-4 Turbo 0.82 0.89 0.85 0.76 0.97 0.85 0.71 0.94 0.81 0.73 0.97 0.83
+GPT-4o 0.78 1.00 0.88 0.67 1.00 0.80 0.57 1.00 0.73 0.55 1.00 0.71
+CodeLlama-7b 0.57 0.92 0.70 0.54 0.94 0.69 0.54 1.00 0.70 0.56 0.94 0.70
+CodeLlama-13b 0.83 0.28 0.42 0.63 0.61 0.62 0.64 0.83 0.72 0.64 0.89 0.74
+Gemini 1.5 Pro 0.85 0.92 0.88 0.67 0.89 0.76 0.67 0.89 0.76 0.61 0.97 0.75
 JavaScript
-GPT-3.5 Turbo 0.74 0.69 0.86 0.77 0.69 0.63 0.93 0.75 0.81 0.85 0.76 0.80 0.72 0.67 0.90 0.76
-GPT-4 Turbo 0.84 0.88 0.79 0.84 0.83 0.77 0.93 0.84 0.78 0.74 0.86 0.79 0.72 0.67 0.90 0.76
-GPT-4o 0.67 0.61 0.97 0.75 0.66 0.59 1.00 0.74 0.53 0.52 1.00 0.68 0.53 0.52 0.97 0.67
-CodeLlama-7b 0.64 0.58 0.97 0.73 0.55 0.53 0.97 0.68 0.52 0.51 1.00 0.67 0.66 0.60 0.90 0.72
-CodeLlama-13b 0.53 0.57 0.28 0.37 0.69 0.72 0.62 0.67 0.66 0.64 0.72 0.68 0.62 0.58 0.86 0.69
-Gemini 1.5 Pro 0.88 0.87 0.90 0.88 0.64 0.60 0.86 0.70 0.71 0.67 0.83 0.74 0.60 0.56 1.00 0.72
+GPT-3.5 Turbo 0.69 0.86 0.77 0.63 0.93 0.75 0.85 0.76 0.80 0.67 0.90 0.76
+GPT-4 Turbo 0.88 0.79 0.84 0.77 0.93 0.84 0.74 0.86 0.79 0.67 0.90 0.76
+GPT-4o 0.61 0.97 0.75 0.59 1.00 0.74 0.52 1.00 0.68 0.52 0.97 0.67
+CodeLlama-7b 0.58 0.97 0.73 0.53 0.97 0.68 0.51 1.00 0.67 0.60 0.90 0.72
+CodeLlama-13b 0.57 0.28 0.37 0.72 0.62 0.67 0.64 0.72 0.68 0.58 0.86 0.69
+Gemini 1.5 Pro 0.87 0.90 0.88 0.60 0.86 0.70 0.67 0.83 0.74 0.56 1.00 0.72
 Java
-GPT-3.5 Turbo 0.60 0.57 0.74 0.65 0.58 0.56 0.76 0.65 0.65 0.69 0.57 0.62 0.58 0.57 0.69 0.62
-GPT-4 Turbo 0.80 0.86 0.71 0.78 0.79 0.76 0.83 0.80 0.82 0.83 0.81 0.82 0.83 0.78 0.93 0.85
-GPT-4o 0.82 0.75 0.98 0.85 0.79 0.70 1.00 0.82 0.68 0.61 1.00 0.76 0.71 0.64 1.00 0.78
-CodeLlama-7b 0.56 0.54 0.88 0.67 0.54 0.52 0.95 0.67 0.54 0.52 1.00 0.68 0.56 0.54 0.76 0.63
-CodeLlama-13b 0.60 0.75 0.29 0.41 0.71 0.78 0.60 0.68 0.68 0.65 0.79 0.71 0.65 0.62 0.79 0.69
-Gemini 1.5 Pro 0.75 0.74 0.76 0.75 0.76 0.70 0.93 0.80 0.77 0.74 0.83 0.79 0.71 0.65 0.95 0.77
-II-D Classification Methods
-Our study examined three classification approaches:
-  • Vulnerability Detection: A simple binary classification method in which the LLM is tasked to categorize an input snippet into either “vulnerable” or “not vulnerable”.
-  • CWE classification (zero-shot): A multi-class classification method in which the LLM is tasked to categorize an input snippet as pertaining to a specific CWE.
+GPT-3.5 Turbo 0.57 0.74 0.65 0.56 0.76 0.65 0.69 0.57 0.62 0.57 0.69 0.62
+GPT-4 Turbo 0.86 0.71 0.78 0.76 0.83 0.80 0.83 0.81 0.82 0.78 0.93 0.85
+GPT-4o 0.75 0.98 0.85 0.70 1.00 0.82 0.61 1.00 0.76 0.64 1.00 0.78
+CodeLlama-7b 0.54 0.88 0.67 0.52 0.95 0.67 0.52 1.00 0.68 0.54 0.76 0.63
+CodeLlama-13b 0.75 0.29 0.41 0.78 0.60 0.68 0.65 0.79 0.71 0.62 0.79 0.69
+Gemini 1.5 Pro 0.74 0.76 0.75 0.70 0.93 0.80 0.74 0.83 0.79 0.65 0.95 0.77
+2.4. Classification Methods
+Our study examined three classification approaches.
+Vulnerability Detection. A simple binary classification method in which the LLM is tasked to categorize an input snippet into either “vulnerable” or “not vulnerable”.
+CWE classification (zero-shot).
+A multi-class classification method in which the LLM is tasked to categorize an input snippet as pertaining to a specific CWE.
 The zero-shot setting provides only a task description in natural language as part of the input prompt.
-  • CWE classification (few-shot): Same multi-class classification method as before, but using a few-shot setting in which the LLM is provided with code examples from the list of the top 25 CWEs [25].
-II-E Prompt and Input Data
+CWE classification (few-shot).
+Same multi-class classification method as before, but using a prompt that provides the top-25 CWE labels (1) to elicit CWE-specific reasoning and map the input snippet to the corresponding CWE category.
+2.5. Prompt and Input Data
 LLM’s prompts were divided into two main types: system prompt, which defines the behavior, and user prompt, which gives instructions regarding a given task.
 In our study, the final prompts combine one system prompt with one user prompt.
-Table III reports all the prompts that were used for our experiments.
+Table 4 reports all the prompts that were used for our experiments.
 We considered two system prompts: VD-Sys1 gives the LLM a detailed role for vulnerability detection.
 In contrast, VD-Sys2 had a more generic role, as described by Zhou et al.
-[37].
-Similarly, we designed two different types of user prompts to detect a vulnerability: VD-User1 presents a concise request for vulnerability detection, whereas VD-User2 provides more descriptive instruction as used by Zhou et al.
-[37].
+(Zhou et al., 2024b).
+Similarly, we designed two different types of user prompts to detect a vulnerability: U1 presents a concise request for vulnerability detection, whereas VD-User2 provides more descriptive instruction as used by Zhou et al.
+(Zhou et al., 2024b).
 For the classification of CWE (RQ2), we introduced two system prompts (CWE-Sys1 and CWE-Sys2) and two user prompts (CWE-UserZ and CWE-UserF).
-CWE-Sys1 is a similar prompt compared to VD-Sys1, which provides a more specific role for CWE classification. Conversely, CWE-Sys2 is the same prompt as VD-Sys2 [37].
+CWE-Sys1 is a similar prompt compared to VD-Sys1, which provides a more specific role for CWE classification.
+Conversely, CWE-Sys2 is the same prompt as VD-Sys2 (Zhou et al., 2024b).
 CWE-UserZ is a user prompt for CWE classification with a zero-shot setting (i.e., without any examples).
@@
 To form the final prompt text, we concatenated an input code snippet with one of our described prompt strategies.
-II-F Procedure and Metrics
-To answer RQ1 and RQ2, we executed each LLM on all code snippets of our datasets, using the prompt strategies described in Section II-E.
+2.6. Procedure and Metrics
+To answer RQ1 and RQ2, we executed each LLM on all code snippets of our datasets, using the prompt strategies described in Section 2.5.
 As evaluation metrics, we employed four metrics, namely accuracy, recall, precision, and $F_{1}$ score, commonly used to evaluate the performance of LLMs in vulnerability detection.
 In our evaluation, valid detections/classifications were treated as the positive class, while detection/classification augmentations were considered the negative class.
-We accounted for CWE hierarchical relationships, considering positive scores for cases where the LLM predicted a code snippet labeled with a specific CWE (e.g., CWE-77) as one of its child CWEs (e.g., CWE-78), according to the MITRE CWE database [38].
+We accounted for CWE hierarchical relationships, considering positive scores for cases where the LLM predicted a code snippet labeled with a specific CWE (e.g., CWE-77) as one of its child CWEs (e.g., CWE-78), according to the MITRE CWE database (MITRE, 2024).
 For CWE classification (RQ2), we evaluated the overall performance by macro-averaging the precision parameter across all classes in the precision, recall, and $F_{1}$ score.
-Overall, in our evaluation, we computed 18,144 predictions (6 LLMs $\times$ 378 code snippets $\times$ 8 prompt configurations), approximately 10 hours of computing time.
-Table V: RQ2: Effectiveness comparison of six LLMs using four prompt configurations for CWE classification. The best scores are indicated in bold.
-LLM / Prompt CWE-Sys1 + CWE-UserZ CWE-Sys2 + CWE-UserZ CWE-Sys1 + CWE-UserF CWE-Sys2 + CWE-UserF
-Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$ Acc. Prec. Rec. $F_{1}$
+Overall, in our evaluation, we computed 18,144 predictions (6 LLMs $\times$ 378 code snippets $\times$ 8 prompt configurations), which accounted for 10 hours of computing time.
+Table 6. RQ2: Effectiveness comparison of six LLMs using four prompt configurations for CWE classification (Precision, Recall, $F_{1}$).
+LLM / Prompt C-Sys1 + UZ C-Sys2 + UZ C-Sys1 + UF C-Sys2 + UF
+Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$ Prec. Rec. $F_{1}$
 All
-GPT-3.5 Turbo 0.45 0.14 0.12 0.12 0.38 0.14 0.12 0.12 0.52 0.34 0.35 0.32 0.44 0.24 0.23 0.21
-GPT-4 Turbo 0.56 0.21 0.18 0.18 0.67 0.30 0.23 0.25 0.61 0.42 0.52 0.44 0.69 0.54 0.58 0.53
-GPT-4o 0.62 0.23 0.19 0.20 0.52 0.22 0.19 0.19 0.67 0.61 0.69 0.61 0.65 0.60 0.74 0.63
-CodeLlama-7b 0.22 0.12 0.09 0.08 0.29 0.08 0.05 0.06 0.11 0.13 0.10 0.08 0.13 0.11 0.08 0.06
-CodeLlama-13b 0.33 0.09 0.08 0.08 0.25 0.08 0.05 0.05 0.23 0.09 0.07 0.07 0.21 0.10 0.09 0.07
-Gemini 1.5 Pro 0.53 0.15 0.13 0.13 0.55 0.16 0.15 0.14 0.66 0.52 0.56 0.51 0.69 0.56 0.52 0.49
+GPT-3.5 Turbo 0.14 0.12 0.12 0.14 0.12 0.12 0.34 0.35 0.32 0.24 0.23 0.21
+GPT-4 Turbo 0.21 0.18 0.18 0.30 0.23 0.25 0.42 0.52 0.44 0.54 0.58 0.53
+GPT-4o 0.23 0.19 0.20 0.22 0.19 0.19 0.61 0.69 0.61 0.60 0.74 0.63
+CodeLlama-7b 0.12 0.09 0.08 0.08 0.05 0.06 0.13 0.10 0.08 0.11 0.08 0.06
+CodeLlama-13b 0.09 0.08 0.08 0.08 0.05 0.05 0.09 0.07 0.07 0.10 0.09 0.07
+Gemini 1.5 Pro 0.15 0.13 0.13 0.16 0.15 0.14 0.52 0.56 0.51 0.56 0.52 0.49
 Python
-GPT-3.5 Turbo 0.39 0.25 0.21 0.22 0.38 0.22 0.20 0.20 0.47 0.27 0.32 0.28 0.46 0.32 0.30 0.29
-GPT-4 Turbo 0.61 0.43 0.42 0.40 0.66 0.42 0.35 0.36 0.51 0.53 0.54 0.48 0.59 0.59 0.55 0.52
-GPT-4o 0.54 0.36 0.26 0.27 0.51 0.35 0.31 0.30 0.53 0.58 0.52 0.51 0.58 0.60 0.65 0.60
-CodeLlama-7b 0.13 0.09 0.07 0.07 0.26 0.16 0.10 0.11 0.01 0.00 0.02 0.01 0.11 0.15 0.14 0.11
-CodeLlama-13b 0.29 0.16 0.15 0.14 0.24 0.10 0.09 0.08 0.21 0.17 0.13 0.11 0.18 0.17 0.18 0.12
-Gemini 1.5 Pro 0.54 0.24 0.21 0.20 0.50 0.25 0.22 0.22 0.67 0.61 0.64 0.58 0.71 0.72 0.62 0.62
+GPT-3.5 Turbo 0.25 0.21 0.22 0.22 0.20 0.20 0.27 0.32 0.28 0.32 0.30 0.29
+GPT-4 Turbo 0.43 0.42 0.40 0.42 0.35 0.36 0.53 0.54 0.48 0.59 0.55 0.52
+GPT-4o 0.36 0.26 0.27 0.35 0.31 0.30 0.58 0.52 0.51 0.60 0.65 0.60
+CodeLlama-7b 0.09 0.07 0.07 0.16 0.10 0.11 0.00 0.02 0.01 0.15 0.14 0.11
+CodeLlama-13b 0.16 0.15 0.14 0.10 0.09 0.08 0.17 0.13 0.11 0.17 0.18 0.12
+Gemini 1.5 Pro 0.24 0.21 0.20 0.25 0.22 0.22 0.61 0.64 0.58 0.72 0.62 0.62
 C
-GPT-3.5 Turbo 0.43 0.21 0.20 0.19 0.34 0.20 0.19 0.19 0.44 0.46 0.43 0.39 0.41 0.32 0.33 0.31
-GPT-4 Turbo 0.60 0.38 0.37 0.36 0.70 0.46 0.42 0.43 0.65 0.56 0.66 0.58 0.70 0.67 0.69 0.66
-GPT-4o 0.61 0.37 0.31 0.32 0.48 0.33 0.28 0.29 0.66 0.64 0.74 0.65 0.60 0.62 0.76 0.64
-CodeLlama-7b 0.22 0.19 0.15 0.15 0.19 0.08 0.06 0.07 0.16 0.12 0.20 0.13 0.15 0.12 0.17 0.12
-CodeLlama-13b 0.31 0.18 0.15 0.15 0.22 0.15 0.12 0.12 0.22 0.14 0.13 0.12 0.24 0.20 0.21 0.18
-Gemini 1.5 Pro 0.40 0.22 0.17 0.18 0.53 0.29 0.26 0.26 0.67 0.57 0.60 0.56 0.70 0.67 0.62 0.61
+GPT-3.5 Turbo 0.21 0.20 0.19 0.20 0.19 0.19 0.46 0.43 0.39 0.32 0.33 0.31
+GPT-4 Turbo 0.38 0.37 0.36 0.46 0.42 0.43 0.56 0.66 0.58 0.67 0.69 0.66
+GPT-4o 0.37 0.31 0.32 0.33 0.28 0.29 0.64 0.74 0.65 0.62 0.76 0.64
+CodeLlama-7b 0.19 0.15 0.15 0.08 0.06 0.07 0.12 0.20 0.13 0.12 0.17 0.12
+CodeLlama-13b 0.18 0.15 0.15 0.15 0.12 0.12 0.14 0.13 0.12 0.20 0.21 0.18
+Gemini 1.5 Pro 0.22 0.17 0.18 0.29 0.26 0.26 0.57 0.60 0.56 0.67 0.62 0.61
 C++
-GPT-3.5 Turbo 0.43 0.21 0.19 0.19 0.36 0.25 0.22 0.22 0.54 0.43 0.46 0.41 0.43 0.32 0.32 0.29
-GPT-4 Turbo 0.53 0.39 0.39 0.37 0.62 0.44 0.42 0.41 0.64 0.59 0.68 0.60 0.68 0.57 0.66 0.57
-GPT-4o 0.65 0.42 0.39 0.39 0.51 0.40 0.38 0.36 0.65 0.61 0.75 0.64 0.60 0.59 0.76 0.62
-CodeLlama-7b 0.33 0.17 0.15 0.14 0.36 0.18 0.12 0.14 0.15 0.17 0.18 0.16 0.11 0.09 0.13 0.09
-CodeLlama-13b 0.29 0.18 0.15 0.15 0.25 0.12 0.11 0.10 0.26 0.22 0.20 0.19 0.24 0.27 0.22 0.19
-Gemini 1.5 Pro 0.58 0.32 0.28 0.29 0.62 0.32 0.30 0.30 0.60 0.45 0.40 0.39 0.64 0.45 0.42 0.41
+GPT-3.5 Turbo 0.21 0.19 0.19 0.25 0.22 0.22 0.43 0.46 0.41 0.32 0.32 0.29
+GPT-4 Turbo 0.39 0.39 0.37 0.44 0.42 0.41 0.59 0.68 0.60 0.57 0.66 0.57
+GPT-4o 0.42 0.39 0.39 0.40 0.38 0.36 0.61 0.75 0.64 0.59 0.76 0.62
+CodeLlama-7b 0.17 0.15 0.14 0.18 0.12 0.14 0.17 0.18 0.16 0.09 0.13 0.09
+CodeLlama-13b 0.18 0.15 0.15 0.12 0.11 0.10 0.22 0.20 0.19 0.27 0.22 0.19
+Gemini 1.5 Pro 0.32 0.28 0.29 0.32 0.30 0.30 0.45 0.40 0.39 0.45 0.42 0.41
 JavaScript
-GPT-3.5 Turbo 0.53 0.25 0.26 0.24 0.47 0.27 0.27 0.25 0.64 0.42 0.50 0.44 0.50 0.41 0.40 0.37
-GPT-4 Turbo 0.52 0.28 0.28 0.26 0.76 0.52 0.52 0.50 0.53 0.52 0.64 0.53 0.79 0.68 0.71 0.66
-GPT-4o 0.69 0.37 0.42 0.38 0.53 0.29 0.29 0.27 0.74 0.59 0.62 0.59 0.69 0.53 0.61 0.55
-CodeLlama-7b 0.21 0.16 0.20 0.17 0.36 0.25 0.25 0.22 0.14 0.13 0.19 0.13 0.14 0.05 0.12 0.06
-CodeLlama-13b 0.41 0.26 0.30 0.26 0.38 0.27 0.27 0.23 0.17 0.08 0.07 0.06 0.14 0.13 0.11 0.09
-Gemini 1.5 Pro 0.66 0.31 0.35 0.33 0.59 0.27 0.28 0.26 0.72 0.60 0.73 0.63 0.76 0.52 0.52 0.49
+GPT-3.5 Turbo 0.25 0.26 0.24 0.27 0.27 0.25 0.42 0.50 0.44 0.41 0.40 0.37
+GPT-4 Turbo 0.28 0.28 0.26 0.52 0.52 0.50 0.52 0.64 0.53 0.68 0.71 0.66
+GPT-4o 0.37 0.42 0.38 0.29 0.29 0.27 0.59 0.62 0.59 0.53 0.61 0.55
+CodeLlama-7b 0.16 0.20 0.17 0.25 0.25 0.22 0.13 0.19 0.13 0.05 0.12 0.06
+CodeLlama-13b 0.26 0.30 0.26 0.27 0.27 0.23 0.08 0.07 0.06 0.13 0.11 0.09
+Gemini 1.5 Pro 0.31 0.35 0.33 0.27 0.28 0.26 0.60 0.73 0.63 0.52 0.52 0.49
 Java
-GPT-3.5 Turbo 0.46 0.19 0.16 0.16 0.39 0.22 0.18 0.19 0.55 0.28 0.31 0.29 0.42 0.22 0.23 0.21
-GPT-4 Turbo 0.51 0.29 0.23 0.25 0.63 0.42 0.35 0.37 0.68 0.49 0.55 0.50 0.69 0.48 0.52 0.48
-GPT-4o 0.62 0.35 0.30 0.31 0.58 0.36 0.32 0.32 0.76 0.59 0.63 0.59 0.77 0.63 0.68 0.64
-CodeLlama-7b 0.21 0.14 0.10 0.10 0.31 0.11 0.11 0.10 0.10 0.07 0.11 0.07 0.14 0.09 0.13 0.08
-CodeLlama-13b 0.36 0.17 0.18 0.17 0.23 0.16 0.16 0.14 0.29 0.14 0.15 0.12 0.24 0.15 0.12 0.12
-Gemini 1.5 Pro 0.51 0.28 0.26 0.26 0.51 0.29 0.26 0.26 0.65 0.59 0.69 0.59 0.65 0.52 0.60 0.52
-III Experimental Results
-III-A RQ1 (Vulnerability detection)
-Table IV (vulnerability detection) presents the effectiveness results across the different prompt setups.
+GPT-3.5 Turbo 0.19 0.16 0.16 0.22 0.18 0.19 0.28 0.31 0.29 0.22 0.23 0.21
+GPT-4 Turbo 0.29 0.23 0.25 0.42 0.35 0.37 0.49 0.55 0.50 0.48 0.52 0.48
+GPT-4o 0.35 0.30 0.31 0.36 0.32 0.32 0.59 0.63 0.59 0.63 0.68 0.64
+CodeLlama-7b 0.14 0.10 0.10 0.11 0.11 0.10 0.07 0.11 0.07 0.09 0.13 0.08
+CodeLlama-13b 0.17 0.18 0.17 0.16 0.16 0.14 0.14 0.15 0.12 0.15 0.12 0.12
+Gemini 1.5 Pro 0.28 0.26 0.26 0.29 0.26 0.26 0.59 0.69 0.59 0.52 0.60 0.52
+3. Experimental Results
+3.1. RQ1 (Vulnerability detection)
+Table 5 (vulnerability detection) presents the effectiveness results across the different prompt setups.
 Overall, GPT-4 Turbo and GPT-4o showed the best binary classification scores.
 The former achieved the highest accuracy and precision (0.81), whereas the latter showed the highest recall (1.00) and $F_{1}$ score (0.83).
-Changing the prompt from VD-Sys1 to VD-Sys2 (i.e., a more general role definition) had a positive impact only for GPT-3.5 Turbo and CodeLlama-13b, proving detrimental for the other LLMs.
-Although VD-Sys2 configurations achieved the highest recall results, they only marginally outperformed VD-Sys1 configurations and came at the cost of reduced precision.
-The best balance between detecting vulnerabilities and minimizing false positives was achieved by the prompt configurations VD-Sys1 + VD-User1 and VD-Sys1 + VD-User2 (the first two macro-columns).
-Among these, VD-Sys1 + VD-User1 was the most effective for minimizing false alarms (GPT-4 Turbo) and detecting security vulnerabilities (GPT-4o).
+Changing the prompt from Sys1 to VD-Sys2 (i.e., a more general role definition) had a positive impact only for GPT-3.5 Turbo and CodeLlama-13b, proving detrimental for the other LLMs.
+Although VD-Sys2 configurations achieved the highest recall results, they only marginally outperformed Sys1 configurations and came at the cost of reduced precision.
+The best balance between detecting vulnerabilities and minimizing false positives was achieved by the prompt configurations Sys1–U1 and Sys1 + VD-User2 (the first two macro-columns).
+Among these, Sys1–U1 was the most effective for minimizing false alarms (GPT-4 Turbo) and detecting vulnerabilities (GPT-4o).
 For Python, GPT-4 Turbo and GPT-4o consistently outperformed other models.
-In particular, both models achieved the highest $F_{1}$ score of 0.80 using the VD-User1 task prompt.
-Although Gemini 1.5 pro delivered the highest accuracy, its $F_{1}$ score did not reach more than 0.80 due to its lower recall than GPT-4 Turbo and GPT-4o across all the prompt settings.
+In particular, both models achieved the highest $F_{1}$ score of 0.80 using the U1 task prompt.
+Although Gemini 1.5 pro delivered the highest accuracy, its $F_{1}$ score did not reach more than 0.80 due to its lower recall than GPT-4 Turbo and GPT-4o.
 In C, GPT-4o demonstrated exceptional recall, achieving perfect scores (1.00) across all prompt configurations.
@@
 This revealed the potential to reliably identify true positive vulnerabilities with a smaller sized LLM.
-For C++, both GPT-4 Turbo and Gemini 1.5 Pro showed a solid performance, both achieving the highest $F_{1}$ score of 0.88 using VD-Sys1 + VD-User1.
+For C++, both GPT-4 Turbo and Gemini 1.5 Pro showed a solid performance, both achieving the highest $F_{1}$ score of 0.88 using Sys1–U1.
 Particularly, GPT-4o surpassed other models in recall (1.0) across all prompt sets, which is similar to C.
-In JavaScript, Gemini 1.5 Pro maintained a balance between precision (0.87) and recall (0.90) in VD-Sys1 + VD-User1; thus, it achieved the highest $F_{1}$ score (0.88).
-Although GPT-4o obtained the highest recall of 1.0 using two prompt sets, CodeLlama-7b, a model more than 10 times smaller in size, also reached a recall of 1.0 (VD-Sys2 + VD-User1).
-Lastly, regarding the result in Java, GPT-4 Turbo and GPT-4o showed the best $F_{1}$ score of 0.85 using VD-Sys1 + VD-User1 used by GPt-4o and VD-Sys2 + VD-User2 used by GPT-4 Turbo.
+In JavaScript, Gemini 1.5 Pro maintained a balance between precision (0.87) and recall (0.90) in Sys1–U1; thus, it achieved the highest $F_{1}$ score (0.88).
+Although GPT-4o obtained the highest recall of 1.0 using two prompt sets, CodeLlama-7b, a model more than 10 times smaller in size, also reached a recall of 1.0 (VD-Sys2 + U1).
+Lastly, regarding the result in Java, GPT-4 Turbo and GPT-4o showed the best $F_{1}$ score of 0.85 using Sys1–U1 used by GPt-4o and VD-Sys2 + VD-User2 used by GPT-4 Turbo.
 While other models did not achieved satisfactory performance, GPT-4 Turbo exhibited the highest precision across all prompt sets.
-In particular, VD-Sys1 + VD-User1, the highest precision of 0.86 (GPT-4 Turbo), was approximately 15% better than the second highest one (GPT-4o).
+In particular, Sys1–U1, the highest precision of 0.86 (GPT-4 Turbo), was approximately 15% better than the second highest one (GPT-4o).
 RQ1 (vulnerability detection): Overall, GPT-4 Turbo and GPT-4o achieved the best vulnerability detection scores across multiple languages ($F_{1}=0.81$ and $F_{1}=0.83$).
@@
 Gemini 1.5 Pro performed the best in JavaScript.
-GPT-4-based models were the most effective LLMs in Java.
-III-B RQ2 (CWE classification)
-Table V (CWE classification) presents the effectiveness results across the different prompt setups.
+GPT-4-based models were most effective in Java.
+3.2. RQ2 (CWE classification)
+Table 6 (CWE classification) presents the effectiveness results across the different prompt setups.
 Unlike the detection task, the scores vary significantly across configurations.
@@
 Similarly, GPT-4 Turbo shows substantial improvements with more than double the recall and $F_{1}$ score between CWE-Sys2 + CWE-UserZ and CWE-Sys2 + CWE-UserF.
-For Python, Gemini 1.5 Pro was the most effective model using CWE-Sys2 + CWE-UserF setting. This exhibited the highest accuracy ($>$ 0.70) and $F_{1}$ (0.62).
+For Python, Gemini 1.5 Pro was the most effective model using CWE-Sys2 + CWE-UserF setting. This exhibited the highest accuracy (>> 0.70) and $F_{1}$ (0.62).
 In C, GPT-4 Turbo obtained balanced precision and recall (0.67 and 0.69), which resulted in the highest $F_{1}$ score among all six LLMs using the CWE-Sys2 + CWE-UserF prompt set.
@@
 Across all languages, the few-shot learning approach (CWE-UserF) consistently yielded better results compared to zero-shot prompts.
-RQ2 (CWE classification): GPT-4o achieved the best vulnerability classification scores across multiple languages ($recall=0.74$ and $F_{1}=0.63$).
-GPT-4 Turbo showed strong performance in C and JavaScript.
-Gemini 1.5 Pro outperformed in Python.
-Few-shot prompts consistently outperformed zero-shot approaches across all models and languages.
-IV CodeGuardian
+RQ2 (CWE classification): GPT-4o achieved the best overall vulnerability classification performance across languages ($recall=0.74$, $F_{1}=0.63$).
+GPT-4 Turbo performed best in C and JavaScript, while Gemini 1.5 Pro outperformed in Python.
+Across all models and languages, few-shot prompting consistently outperformed zero-shot.
+4. User Study
+4.1. CodeGuardian
 To help engineers understand potential vulnerabilities in source code within their development workflow, we developed CodeGuardian, a VSCode extension that features just-in-time LLM-powered vulnerability scanning.
@@
 More specifically, a user can highlight code snippets in editor, and CodeGuardian provides instant vulnerability analysis of the highlighted code.
-This feature helps developers identify and address security flaws at a snippet level.
+This feature helps developers identify security flaws at a snippet level.
 Multi-snippet or multi-file analyses are not supported.
-Our study identified GPT-4 Turbo and GPT-4o as the best models for CWE detection and classification (see Section III).
+Our study identified GPT-4 Turbo and GPT-4o as the best models for CWE detection and classification (see Section 3).
 Consequently, CodeGuardian provides users with the option to select these models, using Azure OpenAI Studio APIs in the back end.
@@
 After the vulnerability analysis, the results are shown on the extension panel.
+Additional details of CodeGuardian are available in our replication package (rep-package, 2025).
 The vulnerability analysis in CodeGuardian comprises two stages: first, a vulnerability detection phase, to determine whether the code is vulnerable or non-vulnerable, and second, a CWE classification phase, employing a few-shot setting for vulnerable code.
 Finally, the output of CodeGuardian is a detailed analysis of the CWE classification results.
-V User Study
+4.2. Experiment Design
 We evaluated CodeGuardian by a dedicated research question, with the main objective of providing a qualitative assessment of LLMs for security vulnerability detection.
@@
 To address RQ3, we performed a user study within an industrial setting, which we describe next.
-V-A Experiment Design
 Participants and Grouping.
@@
 We used a 5-point Likert scale for the skills and knowledge questions and a nominal scale for the rest of the questions.
-Participants were divided into two groups based on their responses: an experimental group (using CodeGuardian) and a control group (using only Internet search).
-Each group consists of 11 participants.
-We mainly referred to each participant’s years of experience in software development and cybersecurity to allocate them into equally distributed groups.
-Challenges Design.
-All participants were required to perform a set of challenges during the experiment, which was identifying possible software vulnerabilities.
-In this experiment, a total of six challenges across five programming languages were presented to both groups.
-Table VI shows all the code snippets used for this user study, each snippet contains a specific vulnerability.
-While five snippets were vulnerable, one of the Python codes was used as a Placebo (i.e., not vulnerable).
-Table VI: An overview of the six challenges of our user study.
+Table 7. An overview of the six challenges of our user study.
 No Language Vulnerability
@@
 6 Python CWE-79: Cross-site Scripting
-Independent Variable. The independent variable was the type of tool used by the participants: CodeGuardian (LLM-based) and Internet search/own knowledge.
+Participants were divided into two groups based on their responses: an experimental group (using CodeGuardian) and a control group (using only Internet search).
+Each group consists of 11 participants.
+By balancing expertise in software development and cybersecurity, we ensured that the experimental and control groups had comparable backgrounds.
+We mainly referred to each participant’s years of experience in software development and cybersecurity to allocate them into equally distributed groups.
+Lastly, none of the participants had any experience with CodeGuardian.
+Challenges Design.
+All participants were required to perform a set of challenges during the experiment, which was identifying possible software vulnerabilities.
+In this experiment, a total of six challenges across five programming languages were presented to both groups.
+Table 7 shows all the code snippets used for this user study, each snippet contains a specific vulnerability.
+While five snippets were vulnerable, one of the Python codes was used as a Placebo (i.e., not vulnerable).
+Independent Variable.
+The independent variable was the type of tool used by the participants: CodeGuardian (LLM-based) and Internet search/own knowledge.
+More specifically, Level 1 expects participants to use CodeGuardian to identify vulnerabilities, and Level 2 expects them to use their knowledge.
 Dependent variables. The experiment for this user study has two dependent variables: (1) challenge completion time and (2) challenge completion accuracy.
@@
 This workstation was an EC2 instance running Debian 12 OS, accessible to all participants via a browser-based VNC.
-Qualitative Analysis.
-The participants in the experimental group were asked to answer a post-questionnaire for the qualitative analysis of CodeGuardian.
-The questions focused on assessing the user experience and effectiveness of our plugin.
 Data Analysis.
-To compare dependent variables across the experimental group and the control group, we analyzed the results using the Mann-Whitney U test [39] (with $\alpha=0.05$) and the magnitude of the differences using the rank-biserial correlation to calculate the effect size [40].
-We selected this non-parametric test as completion time and accuracy data were not normally distributed.
-V-B RQ3 (usefulness)
-Figure 1: Challenge completion accuracy between the experimental group (green) and control group (yellow).
-V-C Completion Accuracy
+To compare dependent variables across the experimental group and the control group, we analyzed the results using the Mann-Whitney U test (Mann and Whitney, 1947) (with $\alpha=0.05$) and the magnitude of the differences using the rank-biserial correlation to calculate the effect size.
+We selected this non-parametric test as the completion time and accuracy data were not normally distributed.
+4.3. RQ3 (usefulness)
+Figure 1. Challenge completion accuracy between the experimental group (green) and control group (yellow).
+4.3.1. Completion Accuracy
 Accuracy was calculated based on the amount of correct answers out of all six challenges.
@@
 We used a one-sided research hypothesis since we assumed that the experimental group had a higher percentage of correct answers than the control group.
-Therefore, the null and alternative hypotheses were defined as follows:
-  • $H_{0}$: The two groups are the same in terms of the total challenge completion accuracy.
-  • $H_{1}$: Total challenge completion accuracy in the experiment group was higher than in the control group.
+Therefore, the null and alternative hypotheses were defined as follows: $H_{0}$: The two groups are the same in terms of the total challenge completion accuracy; $H_{1}$: Total challenge completion accuracy in the experiment group was higher than in the control group.
 The experimental group achieved a median accuracy of 1.0 (std = 0.08), while the control group had a median accuracy of 0.33 (std = 0.25).
@@
 This indicates that the usage of CodeGuardian had a strong positive effect on improving the number of correct answers among participants.
-V-D Completion Time
+Figure 2. Challenge completion time between the experimental group (green) and control group (yellow).
+4.3.2. Completion Time
 We measured the total challenge completion time (minutes:seconds) and compared the results statistically between the experimental and control groups.
 A one-sided research hypothesis was used for this test, as we hypothesized that the experimental group would complete all the challenges faster than the control group.
-Hence, the null and alternative hypotheses were defined as follows:
-  • $H_{0}$: The two groups are the same in terms of the total challenge completion time.
-  • $H_{1}$: Total challenge completion time in the experiment group is less than in the control group.
+Hence, the null and alternative hypotheses were defined as follows: $H_{0}$: The two groups are the same in terms of the total challenge completion time; $H_{1}$: Total challenge completion time in the experiment group is less than in the control group.
 Figure 2 depicts the boxplots of completion time over all the challenges.
@@
 Therefore, we reject the null hypothesis and conclude that the experimental group completed the challenges faster than the control group.
-Figure 2: Challenge completion time between the experimental group (green) and control group (yellow).
-Furthermore, Figure 3 illustrates the completion time per challenge.
-The median completion time in the experimental group was less than that of the control group in all challenges.
-Moreover, challenge 5 had the largest difference in performance (completion time) between the experimental and control groups by approximately 3.5 minutes.
-In contrast, challenge 2 in the experimental group showed a slight median difference compared to the control group($M_{exp}$=2:14 and $M_{ctl}$=3:20).
-Additionally, the lowest observed standard deviation for the experimental group was 46 seconds (Challenge 4), whereas the highest was 1 minute and 13 seconds (Challenge 1).
-Similar to the experimental group, the lowest observed median and standard deviation for the control group were 3:20 (Challenge 2) and 50 seconds (Challenge 6), and the highest median and standard deviation were 5 minutes (Challenge 1 and 6) and 1 minute and 34 seconds (Challenge 1).
-Although there were a few outliers in both experimental and control groups, the gaps between the lowest and highest standard deviations were relatively similar.
-Additionally, the experimental group completed the challenges on average by approximately 47% (Challenge 1), 44% (challenge 2), 58% (challenge 3), 59% (challenge 4), 65% (challenge 5), and 55% (challenge 6) faster, compared to the control group.
-Figure 3: Comparison of challenge completion time per challenge between the experimental group (green) and control group (yellow).
-RQ3 (usefulness): Both statistical tests for the challenge completion time and accuracy showed statistically significant improvements in the experimental group compared to the control group.
-CodeGuardian successfully increased the speed and accuracy of completing the challenges with large effect size.
-VI Threats to Validity
-VI-1 Internal validity
-We compared all LLMs under identical experimental settings and on the same benchmark.
-Our implementation of the testing scripts to evaluate the scores was tested thoroughly.
-Moreover, we evaluated the LLMs with a limited set of prompts.
-Our user study involved 22 participants, a reasonable number due to selection constraints (i.e., prior industrial knowledge, experience in software development, and cybersecurity).
-Although the sample size was small, it achieved the necessary statistical test power.
-The varying levels of software development and cybersecurity expertise among our participants, including their familiarity with programming languages and VSCode, could have influenced our results.
-To mitigate this threat, we collected relevant background information through a pre-questionnaire and divided the participants into two groups, striving for an equal distribution of expertise.
-Also, we used the same environment for all participants (a workstation with Debian 12 OS and a connection speed of around 700 Mbps).
-However, the response time of the GPT-4 Turbo API was occasionally slower, especially during periods of high usage due to the limited number of regions where the model is hosted on Azure OpenAI Studio.
-The slower response times could potentially impact user experience and the seamless integration of CodeGuardian into the development workflow.
-These issues were reflected in the users’ responses to the post-questionnaire.
-However, it is important to note that this is a limitation of the current hosting infrastructure on Azure OpenAI Studio.
-VI-2 External validity
-The limited number of LLMs in our evaluation poses a threat in terms of the generalizability of our results, especially with the availability of newer models.
-Our results may not generalize or generalize differently when considering other LLMs.
-To mitigate this issue, we selected the best models available as of 05.2024.
-Likewise, our paper only involved five programming languages (C, C++, Python, Java, and JavaScript).
-This limited scope might not cover the diversity of programming environments and their associated vulnerabilities.
-Furthermore, we used a relatively small number of code challenges in our user study.
-This was due to the limited availability of our participants, most of whom were industry professionals with constrained time.
-As a result, we included only one snippet per language, with an additional placebo snippet for Python.
-While this approach provided valuable insights into the effectiveness of CodeGuardian for vulnerability analysis, it may not fully represent all vulnerabilities and coding scenarios encountered in real-world projects.
-VII Discussion
-VII-A Model Selection
-GPT-4 Turbo often exhibited higher effectiveness compared to the other large language models (see Section III).
-Although GPT-4o is the successor to GPT-4 Turbo, it did not consistently outperform its predecessor.
-GPT-4o achieved higher recall across all prompt setups, indicating it could identify more true positives.
-However, overall, GPT-4 Turbo excelled in three specific prompt setups (VD-Sys1 + VD-User2, VD-Sys2 + VD-User1, and VD-Sys2 + VD-User) primarily due to GPT-4o’s lower precision.
-This highlights a crucial aspect of using LLMs for vulnerability detection: newer versions or successors do not always outperform older models.
-Therefore, model selection should be based on specific security priorities.
-For instance, GPT-4o is preferable for detecting all potential vulnerabilities (i.e., reducing false negatives), whereas GPT-4 Turbo is better for minimizing false alarms.
-This underscores the importance of selecting the right model based on the particular requirements.
-VII-B Model Performance in Each Language
-Our study revealed significant variations in LLM performance across different programming languages for both vulnerability detection and CWE classification tasks.
-GPT-4-based models mostly outperformed other models across all languages.
-However, the strong performance of Gemini 1.5 Pro in JavaScript (vulnerability detection) and Python (CWE classification) indicates that different models may have strengths in particular languages or types of code analysis.
-This underscores the importance of considering language-specific performance when we select for secure code review.
-VII-C Complexity of CWE Classification using LLM
-Compared to vulnerability detection, our experiments reveal significant challenges in CWE classification, with performance varying across different CWE classes and model sizes.
-Larger models like GPT-4o, GPT-4 Turbo, and Gemini 1.5 Pro consistently outperformed smaller ones, particularly when employing few-shot learning techniques.
-For instance, we observed that recall rates of GPT-4o approximately triple using the few-shot setting.
-This improvement suggests that providing relevant examples helps models better understand the classification task.
-However, using few-shot learning does not help for smaller models, such as CodeLlama models due to the limited parameter size restricting their ability to effectively utilize the few-shot examples.
-VII-D User Study
-In our user study, the experimental group consistently completed challenges faster than the control group.
-Specifically, our placebo code analysis (Challenge 5) showed that most participants who used CodeGuardian were able to answer in under approximately two minutes.
-However, the control group’s participants took, on average, almost all of the allotted time (5 minutes).
-Thus, CodeGuardian helped participants quickly identify whether the code was vulnerable or not.
-We believe that this effect is because of the two-stage processing of CodeGuardian (see Section IV).
-Furthermore, in Challenge 4, the most considerable reduction in completion time was found when participants used CodeGuardian.
-Analysis of our pre-questionnaire revealed that the participants were less skilled in JavaScript than in popular programming languages, such as Java or Python.
-Hence, we assume that, although some participants are not familiar with certain languages, CodeGuardian enables developers to analyze code and identify potential vulnerabilities regardless of prior knowledge.
-In addition, accuracy was notably higher in the experimental group than in the control group.
-Evidence suggests that seamless, instant feedback and automated vulnerability analysis by CodeGuardian not only accelerate the analysis process but also correct vulnerability identification.
-Despite the majority of participants having limited knowledge of cybersecurity, according to the pre-questionnaire, the experimental group had an accuracy of more than 0.90 on average.
-In contrast, the accuracy of the control group was less than 0.50.
-This indicates that CodeGuardian effectively supports developers in identifying potential software vulnerabilities without having prior cybersecurity knowledge.
-However, our tool requires strict usage guidelines for junior developers, who might overly depend on their outputs without proper verification.
-Indeed, LLMs are prone to hallucinations, where the output appears plausible but is largely incorrect.
-We believe CodeGuardian could have significant implications for secure software development practices.
-Reducing the time and effort required to identify vulnerabilities allows developers to conduct security checks more seamlessly within their workflow, potentially leading to more secure code being produced at a faster rate.
-The feedback from the experimental group indicated that the majority of participants ($>$ 70%) were satisfied with the speed of the vulnerability scanning process.
-Furthermore, users particularly appreciated how CodeGuardian effectively checked potential vulnerabilities automatically.
-VIII Related Work
-VIII-A Conventional Methods and Tools
-Conventional vulnerability analysis primarily relies on static application security testing (SAST) and dynamic application security testing (DAST).
-SAST examines source code without execution and focuses on syntactic and semantic checks.
-DAST evaluates applications during runtime, mimicking attacker behavior [41].
-However, SAST typically generates a high number of false positives [42], and DAST provides high false negatives, which may miss many vulnerabilities [43].
-In constrast, in this paper we focus on LLM-based approaches to vulnerability detection.
-VIII-B ML and DL in Vulnerability Detection
-Recently, vulnerability detection systems have become more adapted to learning from data to predict and identify vulnerabilities with higher accuracy [16].
-DL approaches, particularly utilizing Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks, have further improved detection accuracy without requiring feature engineering by hand [15, 44].
-Despite the advancement brought by DL, these models treat code as linear sequences, potentially overseeing complex semantic relationships as well as long-range dependencies [45].
-These approaches are challenged by the development of pre-trained LLMs.
-Unlike DL approaches, LLMs such as the ones used in this paper require no explicit training, while achieving high scores in vulnerability detection.
-VIII-C LLMs in Vulnerability Detection and Classification
-Studies have shown that fine-tuned models like BERT and GPT outperform traditional DL models in vulnerability detection [10, 9].
-However, this approach requires substantial computational resources and large, well-annotated datasets.
-In contrast, prompt engineering leverages pre-trained LLMs without any parameter update, offering a more resource-efficient approach.
-Several studies have demonstrated the effectiveness of prompt engineering, such as few-shot learning and chain-of-thought in vulnerability detection and classification [46, 47, 48].
-Specifically, Zhou et al.
-[37] showed that GPT-4, using prompt engineering, outperformed fine-tuned CodeBERT.
-Although prompt engineering does not require a large dataset, the effectiveness of LLM in vulnerability detection and classification heavily relies on the quality of the dataset used for evaluation.
-Recent studies highlighted data quality issues in existing datasets, such as poor data quality, low label accuracy, and high data duplication rates [26, 27], causing misleading results when measuring the effectiveness of LLMs.
-Since this study focused on optimizing input prompts and inference methods without fine-tuning, it was important to ensure the use of a high-quality dataset.
-Moreover, to the best of our knowledge, existing works are limited in terms of language diversity, which is a characterizing element of our work.
-VIII-D Tools Leveraging LLMs for Vulnerability Analysis
-Although ChatGPT is not specifically designed to analyze and detect vulnerabilities, it has shown promising results [49, 14].
-However, it is challenging to integrate ChatGPT seamlessly into the development workflow.
-To address this limitation, researchers have also developed IDE-integrated solutions [50, 51].
-Yet, current tools rely on less advanced LLMs, potentially limiting their effectiveness compared to state-of-the-art LLMs like GPT-4.
-To this aim, we developed CodeGuardian as a widely accessible VSCode extension that makes state of the art LLM models accessible.
-Our user study with industrial participants suggested the benefits and perils of a more pervasive adoption of LLMs for vulnerability detection beyond purely academic studies, a novel contribution to the state of the art.
-IX Conclusions and Future Work
-This paper presented a comprehensive study on the effectiveness of large language models (LLMs) for secure code assessment across multiple programming languages.
-We evaluated six state-of-the-art LLMs on vulnerability detection and CWE classification tasks using our developed multi-language vulnerability dataset.
-Our results demonstrate that GPT-4 Turbo and GPT-4o outperform other LLMs.
-Although the GPT-4-based models often achieved the highest effectiveness in both tasks and different programming languages, Gemini 1.5 Pro showed strong performance in JavaScript and Python.
-Furthermore, we found that few-shot learning techniques significantly improve performance, especially for CWE classification.
-To bridge the gap between research findings and practical applications, we developed CodeGuardian integrating LLM-powered vulnerability analysis into developers’ workflows.
-Our user study with 22 industry-experienced developers showed that CodeGuardian significantly improves both the speed and accuracy of vulnerability detection, with participants completing tasks 66% faster and achieving 203% higher accuracy compared to traditional methods.
-For future work, we should focus on expanding our manually annotated dataset to improve robustness and generalizability.
-Exploring advanced prompt engineering methods such as chain-of-thought and Retrieval Augmented Generation (RAG) could further enhance LLM performance.
-Additionally, improving CodeGuardian’s user interface and incorporating user feedback will make it a more effective tool for secure software development.
+RQ3 (usefulness): CodeGuardian significantly improved speed and accuracy, with large effect sizes.
+5. Threats to Validity
+Internal validity.
+All LLMs were compared under identical settings and on the same benchmark.
+Our testing scripts were validated, but the limited prompt set may affect generalizability.
+We believe the selected models are representative.
+The user study involved 22 participants with prior industrial experience in software development and cybersecurity.
+Although the sample size was small, it provided sufficient statistical power.
+Differences in participants’ expertise (e.g., programming languages, VSCode) may have influenced results.
+To mitigate this, we collected background data, balanced expertise across two groups, and ensured identical environments (Debian 12 workstation, 700 Mbps connection).
+Response times of GPT-4 Turbo occasionally varied due to Azure hosting limitations, which was reflected in post-questionnaire feedback.
+External validity.
+The evaluation covered a limited set of LLMs and programming languages (C, C++, Python, Java, JavaScript).
+Results may differ for other models or languages.
+To reduce this threat, we used state-of-the-art models as of 05.2024.
+The user study also included a small number of code challenges (one per language plus a Python placebo), constrained by participants’ availability.
+While useful for evaluating CodeGuardian, this may not represent the full spectrum of vulnerabilities and real-world scenarios (Dozono et al., 2025).
+6. Discussion
+Model Selection.
+Although GPT-4 Turbo is the predecessor of GPT-4o, it often outperforms other models.
+While GPT-4o achieved higher recall across all prompt setups, its lower precision caused it to underperform in some cases.
+This suggests that newer LLM versions do not always surpass earlier ones.
+Model selection should thus reflect security priorities: GPT-4o is preferable when minimizing false negatives, whereas GPT-4 Turbo better reduces false positives.
+Performance Across Languages.
+LLM performance varied significantly by programming language.
+GPT-4 models generally achieved the best results, but Gemini 1.5 Pro showed strong performance in JavaScript (vulnerability detection) and Python (CWE classification).
+This highlights the importance of considering language-specific performance when using LLMs for secure code review.
+Vulnerability Type Analysis.
+LLMs were especially effective in detecting injection-related vulnerabilities.
+Detection rates for CWE-89 (SQL Injection) and CWE-79 (Cross-site Scripting) exceeded 81% across languages.
+This strength is particularly relevant for web application security.
+The user study further confirmed this: in Challenge 6 (CWE-79), participants using CodeGuardian completed the task 55% faster than the control group.
+CWE Classification Complexity.
+CWE classification proved more difficult than vulnerability detection.
+Larger models (GPT-4o, GPT-4 Turbo, Gemini 1.5 Pro) clearly outperformed smaller ones, especially with few-shot prompts, which tripled GPT-4o’s recall.
+In contrast, few-shot learning did not improve CodeLlama models, suggesting that smaller models cannot effectively leverage examples due to limited capacity.
+User Study.
+The experimental group consistently completed tasks faster and with higher accuracy.
+For instance, in the placebo task (Challenge 5), most participants using CodeGuardian answered in under two minutes, compared to nearly five minutes for the control group.
+Accuracy gains were equally notable: the experimental group achieved over 0.90 on average, versus less than 0.50 in the control group.
+These results show that CodeGuardian accelerates analysis and improves detection, even for developers with limited security expertise.
+However, strict usage guidelines are needed for junior developers, as LLM outputs remain prone to hallucinations.
+7. Conclusions and Future Work
+We presented a comprehensive study on the effectiveness of LLMs for secure code assessment across multiple programming languages.
+We evaluated six state-of-the-art LLMs on vulnerability detection and CWE classification using a newly developed multi-language dataset.
+Our results show that GPT-4 Turbo and GPT-4o generally achieved the best overall performance, while Gemini 1.5 Pro performed strongly in JavaScript and Python.
+To bridge research and practice, we introduced CodeGuardian, a VSCode extension that integrates LLM-powered vulnerability analysis into developers’ workflows.
+In a user study with 22 industry professionals, CodeGuardian improved task completion speed by 66% and accuracy by 203% compared to traditional tools.
+Future work will focus on expanding and refining the annotated dataset to enhance robustness and generalizability, exploring advanced prompting techniques such as chain-of-thought and retrieval-augmented generation (RAG), other foundation models (Gao et al., 2025; Baresi et al., 2025; Maryam et al., 2025; Merabishvili et al., 2026; Weißl et al., 2025) and tasks such as natural language conversations (Giebisch et al., 2025; Sorokin et al., 2026; Habicht et al., 2025), and improving CodeGuardian usability through interface enhancements and developer feedback.